### v2

# Using machine learning to measure evidence of students’ sensemaking in physics courses

Abstract
In the education system, problem-solving correctness is often inappropriately conflated with student learning. Advances in both Physics Education Research (PER) and Machine Learning (ML) provide the initial tools to develop a more meaningful and efficient measurement scheme for whether physics students are engaging in sensemaking: a learning process of figuring out the “how” and “why” for a particular phenomena. In this work, we contribute such a measurement scheme, which quantifies the evidence of students’ physical sensemaking given their written explanations for their solutions to physics problems. We outline how the proposed human annotation scheme can be automated into a deployable ML model using language encoders and shared probabilistic classifiers. The procedure is scalable for a large number of problems and students. We implement three unique language encoders with logistic regression, and provide a deployability analysis on real student explanations from the 2023 Introduction to Physics course at Tufts University. Furthermore, we compute sensemaking scores for all students, and analyze these measurements alongside their corresponding problem-solving accuracies. We find no linear relationship between these two variables, supporting the hypothesis that one is not a reliable proxy for the other. We discuss how sensemaking scores can be used alongside problem-solving accuracies to provide a more nuanced snapshot of student performance in physics class.
I Introduction
Teaching and instruction in undergraduate physics courses has largely relied on problem-solving as the standard method to measure student performance [1, 2, 3, 4, 5, 6]. Common practice is for “real-time” performance to be measured via multiple-choice or single-solution problems, where canonically correct answers determine the student’s knowledge of the core material. Accuracy scores across assignments and examinations, typically coupled with letter grades, act as signals of progress throughout the course as well as final verdicts of student success.
While engaging in problem-solving is a useful experience for students in a physics classroom, using the problem solution as a measure of student learning assumes a direct correlation that may not always hold. Problem-solving accuracy as a measurand assumes that students will engage in a learning process involving the core material to obtain the problem solution. Often times, there are alternative strategies for obtaining a problem solution such as rote-memorization of the rules or procedures required for solving similar problem types [7]. In this scenario, students would score very high on exams that contain these problem types; however given a previously unseen problem structure where the same core material is to be applied, the students would struggle. Here, a risk of using problem-solving accuracy as the predominant metric is an inflated sense of confidence in both the instructor and the student that the core material has been learned. It could also pose a risk for confounding variables in research studies that aim to investigate how instructional techniques influence student learning [8, 9, 10, 11, 12].
Over the last fifteen years, researchers in STEM Education (Ed.) have focused more closely on modeling a student’s learning process directly - a process typically referred to as sensemaking [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. Different from memorizing facts, procedures, or problem types, sensemaking is a macro-cognitive process that allows one to construct a model of “how” and “why” with respect to the core material that is consistent with prior experiences and integrates new ones [13]. In physics courses, there is an emphasis on physical sensemaking, where the goal is for students to construct models that describe physical phenomena. This is separate from (blended) mathematical sensemaking, which focuses on how students understand equations and symbolic forms as physical concepts across various scientific disciplines [26, 27, 28, 29, 30, 31, 32, 33, 34].
Prior work has primarily focused on characterizing physical sensemaking through numerous qualitative analyses of in-moment social dynamics and individual student interviews [25, 23, 14, 15, 16, 17, 18, 19, 20]. In these scenarios, students were given a problem to solve, but the research objective was focused on modeling the sensemaking process that took place as a problem-solving strategy. These empirical investigations resulted in recognizable qualitative outcomes of the physical sensemaking process. It is thus a natural next step to consider whether these outcomes can be turned into a post-hoc measurement scheme that reliably quantifies evidence of physical sensemaking given student explanations. Crucially, we need such a scheme that scales well with student population size.
In this paper, we take this next step forward and answer three questions: (1) How can we measure evidence of students’ post-hoc physical sensemaking across multiple problems? (2) How can we use machine learning tools to automate this measurement procedure? (3) How can these measurements be used to obtain a more informed picture of student performance in a physics course?
Our main theoretical contribution is a measurement scheme, which quantifies the evidence of students’ physical sensemaking post-hoc by annotating their written explanations for their solutions to physics problems. We outline how this human annotation scheme can be transformed into a deployable machine learning (ML) model using language encoders and shared probabilistic classifiers, freeing up time for a human annotator. We implement three ML model variations with different text encoders, and provide a deployability analysis on real student explanations to four problems from the 2023 Introduction to Physics course at Tufts University. A pipeline based on the BERT language encoder [35] that was fine-tuned for our task achieves the best Area Under the Receiving Operator Characteristic (AUROC) score of on the test set, which contains explanations from students not seen in training data on the same problems as in the training data. Lastly, we compute sensemaking scores for all students, and analyze these measurements alongside their corresponding problem-solving accuracies. The computed Point-Biserial correlation coefficients between the two variables signify that we cannot claim there exists a linear relationship between problem-solving accuracy and evidence of physical sensemaking in the classroom. This demonstrates a need to measure evidence of the sensemaking process explicitly, rather than relying on problem-solving accuracy as a proxy.
Alongside this manuscript, we contribute practical materials that can be used to help other individuals measure physical sensemaking: (1) examples of implementing the measurement scheme on four physics problems with the corresponding human annotated dataset (385 total student explanations) and (2) ready-to-implement code that generates physical sensemaking measurements for new student explanations on the same set of four problems. Both of these materials can be found in the Github Repository “AuSeM” 111Github repo: https://github.com/tufts-ml/AuSeM.git.
II How to Measure Post-Hoc Physical Sensemaking
II.1 Defining Physical Sensemaking
Given a phenomena of interest, physical sensemaking is the macro-cognitive process that constructs the mental model of “how” and “why” for a given physical phenomena by integrating prior and new experiences until consistency is achieved [13]. Inspired by prior works across research fields that have attempted to theoretically model and empirically characterize student sensemaking [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 37] we summarize the key features of this process: disassembling the phenomena into components, identifying characteristics (i.e. feature observations), and considering both descriptive (i.e. associations) and mechanistic (i.e. cause and effect) relationships. During the process, a student may test various hypotheses to “figure things out.” When someone has sensemade, they have formed an organized mental model of these components, characteristics, and relationships that is self-consistent. The phenomena can then be explained with natural and mathematical language. These prior works, which we discuss in greater detail in Sec. II.4, provide the foundational starting point for a quantitative post-hoc measurement scheme. Let us consider a problem which contains a physical phenomena to be sensemade as a part of the solution strategy. Provided a student’s text explanation of how they solved the problem, a human annotator can ask themselves the following questions:
1.
Does the student identify the components that are relevant to the phenomena?
2.
Does the student identify the characteristics of the components that are relevant to the phenomena?
3.
Does the student identify the descriptive and mechanistic relationships that are relevant the phenomena?
These initial questions are general and can be further divided into criteria questions that capture more specific outcomes of the sensemaking process. Thus, we propose 8 domains that represent what we expect to observe in the student’s explanation as a result of physical sensemaking.
1. Objects. The central entities that interact to form the physical phenomena in a single hierarchical level. Each object object interaction is relevant to the phenomena. Examples: “car”, “box”, “ball”, “people”, “particles”, etc.
2. Influences. The non-central environmental entities that interact with objects. Only the influence object interaction is relevant to the phenomena. Examples: “Earth’s gravitational field”, “human interference”, “electromagnetic field”.
3. Properties. The relevant qualities that are confined to the physical boundary of an object. Examples: mass, temperature, volume.
4. Positioning. The spacial states or orientations of an object. Examples: “downwards”, “upwards”, “left”, “right”, “x-direction”, “y-direction”.
5. Movements. The motion states of an object. Examples: “initial velocity”, “slowing down”, “speeding up”.
6. Interactions. The exchanges between all object object and influence object pairs. Examples: “gravitational force”, “human pushing”, “electromagnetic force”.
7. Descriptive Relationships. The associations between properties, positioning, and movements of an object. Examples: “a change in position is a velocity of an object”, “a change in temperature is associated with a change in velocity of a particle”.
8. Mechanistic Relationships. The changes that result from object object interactions and influence object interactions. Examples: “constant acceleration of a ball caused by Earth’s gravitational field exerting a downward net force”; “acceleration of a book initially at rest caused by a human push”.
Provided a student’s text explanation as to how they arrived at a particular solution for a problem, the annotator can ask themselves whether the student identifies criteria in each of the 8 domains. For example, see Figure 1. This is a student explanation in response to Problem 4 [38] below:
Example: Problem 4 1. The net force on the box is in the positive x direction. Which of the following statements best describes the motion of the box? (a) Its acceleration is parallel to the x axis. (b) Its velocity is parallel to the x axis. (c) Both its velocity and acceleration are parallel to the x axis. (d) Neither its velocity not its acceleration need to be parallel to the x axis. 2. Briefly explain your answer to the previous question.
The physical phenomena for a student to sensemake in this problem is an object’s motion under a net external influence. Note that our measurement scheme is only made possible by the student’s response to part 2, while part 1 only captures problem-solving correctness. We interpret the response as a student’s expression of their problem-solving strategy, and we use it to measure evidence of their physical sensemaking (as shown in Fig 1).
II.2 Post-Hoc Measurement as Multi-Label Binary Annotation
Figure 1 demonstrates how an individual student explanation can be annotated with respect to 8 domains; however note that not all 8 domains are relevant (in Fig 1 the descriptive relationships domain is not relevant to the particular problem) and some domains occur multiple times (e.g. the object has more than one positioning in the context of the overall problem). As such, the full set of yes/no criteria questions across the 8 domains is problem-specific. Thus, for a set of problems, an annotator can follow the proposed steps:
Step 1. Select problems from a particular subject area (e.g. classical mechanics), where there exists a physical phenomena to be sensemade in each problem as a part of the strategy to solution.
Step 2. For each problem and for each domain , determine the set of yes/no criteria questions that are pertinent to the specific problem. For example, in Problem 4 (II.1), there is only a single object and thus with the criteria question Does the student identify the object?. However, there are two positioning criteria questions that are relevant to sensemaking an object’s positioning under a net external influence: Does the student identify the object positioning before the net applied force? Does the student identify the object positioning after the net applied force? The annotator should end up with the total number of criteria questions per problem as .
Step 3. Given a student’s text response to problem , the annotator will use the yes/no criteria questions to obtain a vector of binary digits representing their answers. The resulting criteria vector for a student response is denoted as . To further compute a sensemaking score one can sum the number of and divide by the length of the vector . This score is to be interpreted as the amount of evidence of physical sensemaking. An annotator can also decide to weight the domains (or specific criteria) differently. In this work, we consider all domains and criteria to equally contribute to the sensemaking score. To finish off our example, the criteria vector for the student explanation in Fig. 1 is with the corresponding sensemaking score . The student achieved such a high score by imagining a scenario where multiple interactions are influencing the boxes’ dynamics and communicating the resulting mechanistic relationship between a net external force and an the boxes’ change in motion. The student received a yes for all criteria except for the one in the property domain. The full list of criteria for Problem 4 is displayed in the Appendix A).
II.3 Theoretical Benefits and Limitations
Utility. This annotation scheme is intentionally designed to measure evidence of physical sensemaking given student explanations. The measurements allow for evidence of sensemaking to be quantitatively compared to other numerical values of interest, and distinguished from correctness. For example, evidence of sensemaking is different from measuring sensemaking correctness (i.e. the student is engaging in the sensemaking process with information that is determined by the authority figure to be canonically correct). The current scheme could be easily modified for this other purpose; the annotator would simply need to choose to annotate for both sensemaking and correctness according to the problem (e.g. the student identifies the correct positioning). However in this work, we stick to sensemaking scores that measure evidence of sensemaking, and leave correctness for future studies.
Moreover, we emphasize that low sensemaking scores measured do not indicate that students cannot sensemake. It simply measures that in the current snapshot, evidence of sensemaking is low. There could be several reasons as to why the evidence is low (i.e. the problem itself didn’t encourage sensemaking, the student applied a memorized fact or procedure, etc.). Put concretely, our annotation scheme does not allow one to make any claims about causality.
Flexibility. Individuals can design or select their own problem sets, and then annotate the student explanations according to the same 8 domains. In proposing such an adaptive scheme, we attempt to balance imposing a systematic procedure with allowing the annotator to have enough freedom to make decisions.
Reproducibility. We provide a step-by-step outline of the annotation procedure, concise definitions of the 8 domains, and annotated examples of 385 student explanations across 4 physics problems of different criteria vector lengths. Our hope is that this content leaves little ambiguity for a new annotator looking to reproduce our results or implement the scheme in their own work. However, we acknowledge that some variability among annotators could exist and be quantified in the future with inter-rater reliability studies [39].
Reducibility. Each criteria vector length is the amount of binary information (bits) required to represent the evidence of a student’s sensemaking with respect to a particular problem. It is indeed possible that the 8 domains and criteria could be reduced such that a fewer number of binary variables are needed. While a reduced model form is not currently known; we speculate that such a form could exist.
Scalability. We acknowledge that manual annotation of a large number of student responses would require many hours of labor, often infeasible in practical settings. For example, the student explanation in Fig 1 took minutes to annotate. For a class of one-hundred students, this extrapolates to annotation hours with respect to a single problem. To resolve this scalability challenge, we show how the procedure can be automated by training an ML classification model on student explanations.
II.4 Sensemaking Related Works
Odden and Russ explicitly define sensemaking as “the dynamic process of building an explanation to resolve a perceived gap or conflict in knowledge.” The authors provide examples of individuals engaging in this process, and discuss how sensemaking can be conceptualized in three ways: a stance or framing (i.e. an individuals’ expectation of themselves in their current situation), a cognitive process built on the theory of knowledge integration, and a discourse practice related to argumentation. Each conception addresses what sensemaking is with respect to a different system level of study. Cognitive theories are based on the mental components and interactions happening during sensemaking, whereas framing can be considered one hierarchal level up as a latent description of a student’s current expectations in their environment when sensemaking is taking place. Lastly, discourse analysis is based on studying the content within and form of conversational dialogue between students.
Qualitative analysis has shaped our understanding of how sensemaking begins, functions, and ends for students. Odden and Russ [15] demonstrate how vexing questions stimulate the sensemaking process and can sustain it for longer durations of time than typically observed. Benedict-Chambers et al. [40] provides four specific types of questions focused on guiding students towards sensemaking including explication, explanation, science concept, and science practice questions. To characterize the explicit features of in-moment sensemaking, Odden and Russ [23] propose that the process be modeled as an epistemic game. This model introduces sensemaking as a three step procedure: 1. noticing a gap or inconsistency, 2. generating a mechanistic explanation, and 3. resolving the gap or inconsistency. The authors discuss a precursor to sensemaking, which includes assembling a knowledge framework by initially discussing ideas socially. This brainstorming step is necessary for one to eventually recognize a gap in their own knowledge system. During the game, the students can make moves (e.g. refining a question, appealing to reality, building an analogy, etc.) and if certain constraints are violated, students exit the game play. Three insightful constraints are discussed: 1. sensemaking is a discussion between individuals rather than an imbalanced monologue, 2. sensemaking occurs when there is a phenomena to sensemake that has not already been sensemade, and 3. sensemaking contains self-generated definitions of ideas alongside formal definitions. Kapon [20] puts forth a metric for how explanations generated during the sensemaking process are implicitly evaluated by individuals - based on intuitive knowledge, mechanism, and framing. More specifically, students evaluate explanations as “good enough” based on whether it aligns with their phenomenological experiences, is mechanistic rather than teleological, and satisfies their immediate expectations of their own understanding. Odden [19] explores how conceptual blends support sensemaking through a qualitative case study of students using concepts from kinematics to understand the mechanisms that underlie the phenomena in electric circuits. The author proposes the approach of generalizing conceptual ideas from one sub-area of physics to another as a mechanism that allows students to make connections in the sensemaking process. Furthermore, Sirnookar et al. [18] explores the role of modeling in sensemaking and analyzes the interactions of these two processes as two students are interviewed while solving a physics problem. The authors find that barriers in modeling can inhibit students’ sustained sensemaking.
In pursuit of the longer-term impacts of sensemaking, Radoff et al. [14] provides a qualitative account of a student engaging in sensemaking throughout an academic semester. Interviews and coursework demonstrated that experimentation with sensemaking caused more positive associations with learning, taking on challenges, and constructive confusion. In Kuo et al. [31], the authors show through a controlled classroom experiment that students who are instructed with mathematical sensemaking stimulation are prone to obtaining solutions to physics problems that are more accurate and insightful than those who did not receive this type of explicit instruction.
The sensemaking literature is fragmented across several scientific disciplines and is also typically broken into categories of scientific and mathematical sensemaking, where more recent work has attempted to define a blended model of the two [16, 26, 27]. Kaldaras and Wieman [27] build upon prior works to put forth a cognitive model of blended math-science sensemaking that introduces three levels of sophistication: 1. developing qualitative relationships among relevant variables in mathematical equations that describe a particular phenomenon, 2. developing mathematical relationships among these variables, and 3. explaining how the mathematical operations used in the equation connect back to the physical phenomena. The authors attempt to validate their model by analyzing student responses from interviews, specifically probing the various levels of sophistication in the framework. While the authors found convincing evidence that the levels of sophistication were present across all student responses, it remained inconclusive as to whether there exists an actual sophistication ranking in the levels. As a result, it is still uncertain whether there exists a hierarchy among cognitive connections in math-science sensemaking. This work is directly linked with research that asks: what does it mean to understand a physics equation? [34] and seeks out structure in symbolic expressions that aid in student understanding.
The works above collectively agree on a key characteristic of the sensemaking process: the dynamic generation of mechanistic explanations for a given phenomena. When students have resolved their inconsistency, they have settled on a mechanistic explanation. The emphasis on mechanism in sensemaking is historically tied with research that characterizes student’s mechanistic reasoning in the classroom [17, 25, 41]. Russ et al. is arguably the the most influential contribution made thus far in defining mechanism in the context of STEM Ed., providing a framework for discourse analysis developed from ideas in philosophy of science [25]. In the framework, mechanistic reasoning is characterized as a process that uses experience-based evidence to construct causal relationships among objects describing “how underlying structure and activities can account for observable changes in a system”. The work puts forth a coding scheme of nine categories: 1. describing the target phenomenon they are trying to explain, 2. identifying setup conditions of the spatial and temporal organization of entities that construct the phenomena, 3. identifying the entities as components that construct the phenomena, 4. identifying activities (actions and interactions) of the entities, 5. identifying properties of entities, 6. identifying the spatial organization of entities relative to each other, 7. chaining reasoning strategy for backward or forward inference, 8. analogies, and 9. mental model animation. This framework has been widely adopted by the STEM Ed. community for qualitatively analyzing student reasoning, and we highlight that it overlaps significantly with our 8 domains presented to measure the evidence of sensemaking. There are a few key differences. The first is that we separate entities into objects and influences, where influences interact with objects in such a way that only the influence object effect is relevant to the particular phenomena. This is especially useful when dealing with fields in physics, which are impacted from object interactions, but the effects are oftentimes ignored in classical settings or introductory courses. Our two domains positioning and movement track both the set-up and subsequent conditions of the dynamics of each object. The positioning domain also includes how the objects relate to one another structurally (i.e. object A is on top of object B) when this is relevant for the particular phenomena. As such, we reduce three of the codes (2,6,9) in Russ et al. to a single domain. As we aim to measure post-hoc explanations rather than in-moment discourse, we can code for the chaining and analogy creation process (which can be viewed as sub-sensemaking processes) proposed in Russ et al. by annotating whether the student identifies the relevant mechanistic relationships. In addition, we add a domain for descriptive relationships, which are the relations between the properties of objects - more akin to correlation than causal mechanism. Otherwise, with respect to Russ et al., our properties domain directly maps to code 5 and our interactions domain directly maps to code 4. As previously stated, our work builds off of the numerous qualitative studies discussed above that have identified the features of sensemaking and mechanistic reasoning through analyzing student discourse. We view our main contribution as transforming these features into 8 domains that contain sub-questions with binary outcomes such that we can obtain a post-hoc sensemaking measurement of a student’s explanation with respect to any particular physics problem. Introducing such a binary encoding scheme allows one to quantify evidence of student sensemaking and naturally lends itself to automating the measurement scheme with ML.
III How to Automate Measurements
III.1 Multi-Label Binary Classification Model
ML models have been shown to be capable of replacing labor-intensive encoding of qualitative data with automated procedures [41, 42, 43, 44, 45]. In this work, we only consider models which are open-access, which allows us to train classifiers for our specific task and does not require us to share student data with a 3rd party. Furthermore, we limit our consideration to models which are capable of running on consumer hardware. In our scenario, the task is the same as the human annotator’s: to map student text explanations across independent problems to sensemaking scores .
From text to embeddings. Given a student explanation , we can select an encoder to map all text in the response to a numerical representation . A simple encoder design is counting; we determine a vocabulary as the set of words across all student explanations in the training dataset, with an additional vocabulary entry for unknown words, and record the number of times each one appears in the example. This numerical representation, or embedding , typically called BagofWords, reduces the richness of the text explanation to the pattern of word occurrence [46, 47]. From here on, we will refer to this encoder as BoW.
To capture more complex patterns (e.g relative co-occurrence between all words in the explanation, in-context word occurrence, and word positional associations with grammar rules), one can use a large language model (LLM) encoder [48]. For example, Google’s BERT (Bi-directional Encoder Representations from Transformers) model combines tokenization, self-attention, and a multi-layer perceptron neural network - all pre-trained on English Wikipedia and BookCorpus to generate token embeddings [35]. A token is a set of characters containing a unique ID in BERT’s vocabulary, such that an entire text can be converted to tokens with individual embeddings , which carry information about all other tokens in the sentence. BERT also uses a sentence-level embedding, known as , which can be used as a full-response representation rather than using distinct representations for each token. From here on, we will refer to this encoder as BERTfrozen. Additionally, can be fine-tuned by training BERT’s previously trained weights on new downstream classification tasks [49]. We will refer to this encoder as BERTfinetuned. While requiring more computational resources, this allows for the embedding to capture patterns in the text that are specific to the target prediction task.
Shared classifiers across problems. Given a generated embedding for a student explanation, we want to classify the representation into its correct criteria vector . Thus, we have a multi-label task [50]. If we construct separate binary classification models for each criteria question, then we are required to train independent classifiers. Rather, we propose to work off of an informed assumption: shared domains will have similar word occurrence and grammatical structure across P problems. For example, the phrase “speeding up” should indicate evidence of sensemaking in the movement domain regardless of the problem it appears in , and words that should be classified as objects will share similar positions in the sentence across problems due to object-subject-verb rules. Thus, for problems, we can train 8 independent classifiers. We reduce from a scaling linear in to constant in . Note that we are proposing to train 8 classifiers to obtain annotations for student explanations across problems, where each problem has a different number of criteria questions per domain . Thus, to obtain an annotation for each criteria question within a domain, we add a criteria context embedding to each text embedding for a given student explanation. The updated embedding vector is denoted as . The index specifies the criteria question that we would like for the classifier to label (e.g. Does the student identify the object positioning before applied force (e.g. upwards)?). We note that there are multiple possible ways to encode this context. For example, rather than adding the vectors to obtain the updated embedding, one could concatenate the two vectors together. Additionally, one could embed the entire problem text alongside the specific criteria question. As a starting point, we’ve selected our approach of adding the embeddings together and only including the criteria question. We leave the exploration of alternative methods for future work.
Figure 2 visually illustrates how the model makes a probabilistic prediction answering a specific yes/no criteria question given a student-generated text response. Given embedding , we use a domain-specific logistic regression (LR) classifier to produce the predicted probability . We set , where are the weight vector and bias scalar for domain , and denotes the logistic sigmoid function.
The figure also shows how that prediction impacts the loss used to train the model. For training, we have a dataset student embeddings with human annotated labels (the criteria vectors) . Each represents a single binary label with respect to criteria question of domain for the student explanation . Each is assumed conditionally independent of other labels given and all model parameters.
We minimize the following binary cross entropy loss with an added regularization term over the weights averaged over all explanations, domains, and criteria:
For the precise loss notation utilized in our methods including regularization terms, see Appendix C. The training procedure encourages all 8 classifiers to accurately predict their assigned pieces of the criteria vector for an embedded student explanation .
For some encoder choices, we do not update any encoder parameters during training; only the classifier weights and biases are updated via gradient descent. When fine-tuning BERT, our gradient descent procedure updates encoder parameters alongside classifier weights. Both are explored in our deployability analysis in IV.1.
III.2 Theoretical Benefits and Limitations
Utility. The proposed model maps new student text explanations across a set of problems to sensemaking scores , freeing up time for an annotator. However, the current version of this model is limited to measuring sensemaking scores of student explanations on the same set of problems. We hypothesize that with more labeled training data across new problems, the model proposed here will have enough data to perform well on unseen problem instances in physics. However, the amount of data needed to cross such a threshold is unclear. Regardless, initial efforts to address this limitation should be directed towards collaborative annotation.
Reproducibility. The automation scheme ensures that the same student explanation receives the same annotation. The process is fully deterministic given fixed seeds to control for randomness. Randomness is only present in the initialization of the classifier weights and biases.
Sustainability. As ML research advances, new encoders may be able to improve the overall classification performance. However, many of these models require High Performance Computing (HPC) access and large amounts of energy resources for training [51]. As language modeling in ML improves, it is worth benchmarking new methods while also taking into account the amount of compute resources required to obtain only modestly better performance.
III.3 ML for Education Related Works
Machine learning tools are slowly integrating into STEM Ed. research methods, mostly with the introduction of text classification models [52, 43, 44, 45, 53]. One focus is on in-context learning, where student text responses are directly fed into a chatbot window of a large language model to be labeled [42]. This procedure can be conducted with a few examples (or zero) of student text, opening up an opportunity for non-ML practitioners to seamlessly use computation for data annotation. However, this procedure relies on the assumption that pre-trained language models are useful for specific tasks in education (e.g. classifying correctness, reasoning, sensemaking, etc.). Other works focus on training supervised ML models from scratch for their specific task [44, 45, 41]. For example, Watts et al. [41] uses the mechanistic reasoning approach from Russ et al. to analyze student explanations in organic chemistry with the aid of ML binary classifiers. This work determines convolutional neural networks to outperform other model architectures for this task, but does not provide comparison results to other models. In our work, we provide a model comparison with three different language encoders (BoW, BERTfrozen, and BERTfinetuned), where we also fine-tune a pre-trained large language model on data for our specific task. Furthermore, we use a multi-label scheme that allows us to take advantage of shared structure across problems by providing each classifier with additional context for the problem-specific criteria we care about. As such, our classifiers are able to train on and classify new data across different physics problems without introducing new classification heads.
In response to the increasing amount of supervised ML models in STEM Ed. research, Fussell et al. [54] effectively summarizes and provides a demonstration of the best practices widely used by ML practitioners for validating these ML models. We note that our work utilizes methods advised by this work, including both cross-validation and bootstrapping. Lastly, we note that while our work does not fall under the category of unsupervised learning, these models are a promising avenue of research in STEM Ed. and can be used to cluster student data, detect static latent variables in student text, and analyze the latent dynamics within student conversations [55].
IV Practical Implementation Results
IV.1 ML Model Deployability Analysis
We seek to determine the deployability of the proposed classification models outlined in Sec. III.1. More specifically, our intention is to estimate how these models will perform given explanations from new students that it has never seen before, but on the same set of problems. Below, we outline our implementation and evaluation procedures prior to showcasing final results.
Dataset of student explanations. Our dataset consists of 385 total de-identified student explanations across four problems distributed to students in the first few lectures of the 2023 Tufts University Introduction to Physics course. The retrospective analysis of this deidentified human subjects data was approved by our institutional review board. All problems have the same structure as Problem 4 II.1 shown above and are presented to the students as pre-lecture digital assignments [38]. The same students are asked to solve each problem and explain how they obtained their solution. A few students chose not to respond to a particular problem, which brings our total number of explanations to . The problems capture different general core material for students to sensemake: (1) the descriptors of an object’s motion and dynamics; (2) the model of an object’s motion and dynamics under Earth’s gravitational influence; (3) the model of the relative motion of objects; (4) the model of an object’s motion dynamics under a net external influence. These problems vary in criteria vector size: (see Appendix A for exact problems and their details from [38]). The annotation procedure was carried out by a domain expert: the first author of this work who holds a doctorate in physics.
Model and training specifics. In our analysis, we use the following models BoW, BERTfrozen, and BERTfinetuned. To evaluate performance, we randomly select 15 students who responded to all four questions to use as a held-out test set and use the rest of the data for training and validation (325 samples). To prevent data leakage, we stratify the data by students to ensure that responses from the same student do not end up in both training and test/validation sets.
All models are trained with the Adam optimizer [56] for epochs. We perform a grid search over key model hyperparameters: learning rate and regularization strength ( for classifier weights and for BERT model weights), the full-details of which are available in Appendix D. For BoW and BERTfrozen, we use a 5-fold cross-validation approach for hyperparameter selection, where we select the hyperparameter combination with the best final average AUROC score across folds. AUROC is the computed area under a Receiver Operating Characteristic (ROC) curve defined by the true positive rate vs. the false positive rate of the predicted labels at different classification thresholds. There are two main reasons we believe AUROC to be the most useful metric for our context: (1) It does not require us to choose an arbitrary probability threshold to distinguish our positive predictions (label 1) from our negative predictions (label 0) for each classifier. Instead AUROC sums over all possible thresholds, providing a good summary that is still bounded, AUROC , with increasing values indicating better performance. (2) Due to imbalanced annotations (i.e. some classifiers have more of a positive or negative class rather than a 50-50 split) in our training dataset, AUROC is a safer metric to report than precision or accuracy, as the chance baseline is independent of imbalance. We include an annotation summary of the data in the Appendix B.
The best model for BoW achieved an average AUROC score of across validation folds with the hyper-parameters: . Similarly, for BERTfrozen, the best model achieved an AUROC score of with the hyper-parameters: . For BERTfinetuned, we require a larger amount of computational resources per experiment, so we only used a single training and validation set to run the hyperparameter grid search. Note that we use an L2SP loss [57, 58] when training BERTfinetuned, consisting of two separate weight decay terms such that we can regularize the weights from the BERT encoder separately from the weights of our LR classifiers. This is done to best preserve the prior information learned in in BERT’s original training. We set the batch-size to , which has been shown empirically to work well with deep neural networks [59]. Overall, we run 324 fine-tuning experiments for BERTfinetuned on Tufts’ HPC services. We picked the model with the optimal final AUROC of on the validation set, which had the following hyperparameters: , , and .
Model evaluations. We retrain our selected models on all training and validation data, and evaluate on the held-out test set of randomly selected students (60 student explanations). To approximate uncertainty on unseen student explanations, we bootstrap our test set with replacement to generate individual test sets for evaluating performance [60, 61]. In Figure 3, we show the average AUROC across independent test predictions for each of the 8 domain classifiers and for each model type. We are not able to report the AUROC for the properties domain as the class imbalance is extremely large and no positive examples were included in the test set. As such, we also provide the Area Under the Precision-Recall curve (AUPRC) results in Appendix E for all domain classifiers. This metric emphasizes the precision of the classifier, which informs us as to how well the model predicts the true positive labels. We note that even with AUPRC, there is too much uncertainty to make concrete claims about the deployability of the properties domain classifier. More positive data examples are necessary for improvement. We show the average AUROC across independent test set predictions for each model in Table 1. These values represent the annotation quality of our entire model. In Table 1, we also report the time taken to train the final selected model. To show the significance in results between models, we provide the confidence interval in Table 1 that results from our bootstrapping procedure and taking the AUROC differences between models. These values can be interpreted as the uncertainty intervals in AUROC difference across models, such that as long as zero is outside of the interval, we can be confident that there exists some non-zero difference between our models indicating that one is indeed better than the other. Additionally, we report the confusion matrices in Appendix F. BoW performs surprisingly well with an AUROC score of and trains efficiently in 1 minute. It performs nearly as well for objects as the other two models. This behavior is unsurprising as BoW reduces text to word frequency counts. BERTfrozen performs slightly better with an AUROC score of and takes 7 minutes to train. It performs the best for two of the domains: objects and mechanistic relationships. It is able to reliably outperform BoW on average, as 0 in not inside of the confidence interval between the AUROC classifier difference reported in Table 1. BERTfinetuned achieves an AUROC score of . It takes 160 minutes to train, but we emphasize that this is not a prohibitive cost for a model which only needs to be trained when new student data is available. It performs the best for influences, positioning, movements, interactions, and descriptive relationships. It is able to perform the best overall despite only leading in 5 domains due to the imbalance in the occurrence of each domain. We recommend the BERTfinetuned model as we believe the average increase in performance across domains is worth the nearly 3 hour training time, and the confidence intervals in 1 indicate it is a significant improvement over BERTfrozen. However, if the annotator wants to optimize for a particular domain, they should refer to Figure 3 to decide the best model for their purpose.
| AUROC (95% CI) | Runtime | |
| BoW | 0.839 (0.803, 0.873) | 1 min |
| BERTfrozen | 0.890 (0.864, 0.915) | 7 min |
| BERTfinetuned | 0.916 (0.893, 0.939) | 160 min |
| Avg. (95% CI) | |
| 0.026 (0.005, 0.045) | |
| 0.077 (0.051, 0.104) | |
| 0.051 (0.024, 0.079) |
IV.2 Student Performance Analysis
We provide a quantitative snapshot of student performance in the first few weeks of the 2023 Tufts Intro. to Physics course. More specifically, we compute the sensemaking scores of students from their explanations to four problems (used in Sec. IV.1 and displayed in Appendix Sec. A), and compare these scores alongside problem-solving accuracies. Students reported their multiple-choice answer, such that accuracy here is a binary variable . Our results in Fig 4 show that assessing sensemaking scores provides a more nuanced view of student performance than correctness alone. In problem 1, we observe a high proportion of students solving the problem correctly () with very low sensemaking scores among correct answers (). In problem 2, we also observe high correctness (), but with higher sensemaking scores (an average of ) among correct answers. In these problems we observe students who obtain a correct answer with both high and low evidence of sensemaking. In fact, a closer look at the data reveals that many of these students who arrived at the correct answer without showing evidence of physical sensemaking deployed a learned mathematical rule (i.e. the derivative of velocity is acceleration).
In problems 3 and 4, we observe less correct answers across students ( and respectively); however sensemaking scores remain high among both correct (average sensemaking scores of and respectively) and incorrect ( and respectively) responses. Specifically in problem 4, the mean sensemaking scores for the incorrect group are higher than for the correct group. Recall that our measurement scheme determines evidence of the sensemaking process, not whether students have accurate information in the sensemaking process. Also, sensemaking has been characterized as a non-linear process to organize information that includes integrating ones phenomenological experiences, which often leads to constructive confusion [13, 14, 15, 16, 17]. Investigating the student responses reveals that students with higher sensemaking scores who arrived at the incorrect answer either used incorrect information or were unable to successfully make connections to their personal experiences with everyday physical objects.
| Problem | CI | |
| 1 | -0.166 | (-0.461, 0.145) |
| 2 | 0.053 | (-0.055, 0.176) |
| 3 | 0.404 | (0.227, 0.546) |
| 4 | -0.198 | (-0.391, 0.009) |
We show the Point-Biserial correlation coefficient , known as the binary-continuous form of the Pearson correlation coefficient ), for the student data in Table 2. The coefficient , where values farther from zero signals a stronger linear relationship and the sign indicates the direction. Only results in problem 3 suggest that there exists a positive linear correlation between problem-solving accuracy and sensemaking score; however, the coefficient is closer to zero than one indicating that the relationship strength is weak (). As we cannot assume normality in the underlying distributions over sensemaking scores for each binary group, we circumvent a hypothesis test with a bootstrapping procedure. We sample our data with replacement times and generate an empirical distribution over scores for each sample. We report the confidence interval in Table 2. All values within each interval remain closer to zero than , except for problem 3 for the tail end of the confidence interval. The weak correlation present makes a clear point: problem-solving accuracy is not a useful proxy for measuring evidence of physical sensemaking, nor vice versa. If one were to only look at the accuracy scores, one might believe that students are learning the core material. Sensemaking scores add more nuance to this belief, suggesting that students are engaging in alternative methods to obtaining a problem solution.
There are multiple actions one could take to further analyze these measurements and use them in a practical context, especially at the individual student level. For example, measurements could be used to observe whether evidence of sensemaking for a particular student is declining, stable, or growing across problems in time. To show an example of this, we show the trend for student in Fig. 4. This student shows evidence of stable sensemaking growth, which does not always coincide with correctness. In fact, if one only observed the accuracy score, the student’s performance would appear to be declining. When both of these measures are combined, one observes a more nuanced picture of the student’s progress, which could lead to more informed interventions.
V Conclusion & Future Work
In this work, we contribute a post-hoc measurement scheme to quantify evidence of physical sensemaking via human annotation of student explanations to physics problems. To streamline the lengthy process of annotating each student explanation, we automate the measurement scheme with a trained ML model using language encoders and shared probabilistic classifiers. We implement three model variants with different language encoders and conduct a deployability analysis on real student explanations from Tufts’ 2023 Intro. to Physics course. We show that the ML models have good average performance on data from new students (0.916 AUROC for the best model), but performance is dramatically different across domains. Moreover, we show that across problems, we can now systematically measure evidence of student learning separately from problem-solving accuracy. In fact, we observe no linear correlation between these two variables. Such a result has important implications for both practical instruction and STEM Ed. research. First and foremost it suggests that we need to be more cautious with how we measure learning. It’s not always useful to make assumptions about the process from the outcome. Students obtaining correct answers does not necessarily indicate a student has engaged in physical sensemaking, as there are alternative strategies, oftentimes requiring less time in the short-term, to achieve the same goal. As previous works support, sensemaking can more directly be measured by paying attention to how students solve problems - whether that’s through a post-hoc written explanation or an in-moment oral discussion. The former is where we’ve focused our efforts, but we believe the latter can also be more efficiently executed with ML tools. Beyond this, we perceive two next-stage research questions that lie at the intersection of ML and STEM Ed. fields:
(1) How can this post-hoc measurement scheme be improved? On the ML side, a natural first reaction is more data. We hypothesize that with more labeled training data across new problems, the models proposed here will be able to achieve better performance on the current task and generalize to unseen problems. Given collaborative annotation efforts, such a model could be realized more efficiently and would save all users time in the long run. We invite interested collaborators to get in touch and encourage open sharing of sensemaking data when possible and appropriate. One could also experiment with new state-of-the-art encoders to enhance overall performance [62]. For example, using the Mistral model (open access, open source transformer model in Ref. [63]) as the LLM text encoder, pre-trained with 7 billion parameters and 2 million language tokens. As off-the-shelf language models improve, one might not need as much training data to obtain quality performance. Lastly, our ML model assumes domain independence in the classification task. Perhaps there is a shared structure across domains that could be exploited. As a brief intuitive example, consider the perspective of an annotator. Given that the annotator identifies criteria in the first two domains (objects and influences), it is more probable that the annotator will identify criteria in the other domains. On the STEM Ed. side, we need to better understand how we can design problems that are aligned with the measurement scheme without creating a new performance metric for students. We do not want the annotation criterion to be perceived by students as tasks to complete. This runs the risk of students inventing alternative strategies to effectively mimic sensemaking. An example of this would be pattern recognizing the types of explanations that lead to high sensemaking scores. As shown by our ML model, this isn’t a very difficult task. Ideally, we prompt students to sensemake and our measurement scheme does not disturb their epistemic framing. There exists an open research pursuit in STEM Ed. to understand how problems can be effectively designed and communicated to students such that genuine sensemaking is encouraged [40].
(2) How can we model the dynamics that lead to low vs. high evidence of physical sensemaking in the classroom? Our proposed measurement scheme only provides evidence of a student’s sensemaking at a single point in time. We lack a model as to how students evolve such that specific sensemaking scores are realized. It is of high interest to construct both deterministic and probabilistic ML mechanistic models that capture short and long-term developmental dynamics of student sensemaking in physics classrooms[64, 65].
Acknowledgements.
We are grateful for financial support from the U.S. National Science Foundation via awards GCR and CAREER . Furthermore, we wish to acknowledge Prof. David Hammer for providing us with the dataset necessary to carry out this work, high-level feedback on the draft, and thoughtful discussions regarding STEM Ed. literature. Without him, this work would not have been possible. We wish to acknowledge other members of the Hughes lab research group, namely Ethan Harvey and Patrick Feeney, for thoughtful ML discussions. Additionally, we’d like to thank the baristas at Jaho (Gabriel) and Starbucks (Ken) for supporting us through the final stages of this project.References
- Adams and Wieman [2015] W. K. Adams and C. E. Wieman, Analyzing the many skills involved in solving complex physics problems, American Journal of Physics 83, 459 (2015).
- Lestari et al. [2021] Lestari, S. Syafril, S. Latifah, E. Engkizar, D. Damri, Z. Asril, and N. E. Yaumas, Hybrid learning on problem-solving abiities in physics learning: A literature review, Journal of Physics: Conference Series 1796, 012021 (2021).
- Docktor et al. [2016] J. L. Docktor, J. Dornfeld, E. Frodermann, K. Heller, L. Hsu, K. A. Jackson, A. Mason, Q. X. Ryan, and J. Yang, Assessing student written problem solutions: A problem-solving rubric with application to introductory physics, Phys. Rev. Phys. Educ. Res. 12, 010130 (2016).
- Burkholder et al. [2020] E. W. Burkholder, J. K. Miles, T. J. Layden, K. D. Wang, A. V. Fritz, and C. E. Wieman, Template for teaching and assessment of problem solving in introductory physics, Phys. Rev. Phys. Educ. Res. 16, 010123 (2020).
- Morphew et al. [2018] J. W. Morphew, J. P. Mestre, H.-A. Kang, H.-H. Chang, and G. Fabry, Using computer adaptive testing to assess physics proficiency and improve exam performance in an introductory physics course, Phys. Rev. Phys. Educ. Res. 14, 020110 (2018).
- Lipnevich et al. [2020] A. A. Lipnevich, T. R. Guskey, D. M. Murano, and J. K. S. and, What do grades mean? variation in grading criteria in american college and university courses, Assessment in Education: Principles, Policy & Practice 27, 480 (2020).
- Fakcharoenphol et al. [2011] W. Fakcharoenphol, E. Potter, and T. Stelzer, What students learn when studying physics practice exam problems, Phys. Rev. ST Phys. Educ. Res. 7, 010107 (2011).
- Ben Ouahi et al. [2021] M. Ben Ouahi, M. Ait Hou, A. Bliya, T. Hassouni, and E. M. Al Ibrahmi, The effect of using computer simulation on students’ performance in teaching and learning physics: Are there any gender and area gaps?, Education Research International 2021, 6646017 (2021).
- Didi ş Körhasan and Pospiech [2024] N. Didi ş Körhasan and G. Pospiech, Examination of students’ quantum physics knowledge construction with peers and metacognitive evaluations, Phys. Rev. Phys. Educ. Res. 20, 020153 (2024).
- Laukenmann et al. [2003] M. Laukenmann, M. Bleicher, S. Fuß, M. Gläser-Zikuda, P. Mayring, and C. von Rhöneck, An investigation of the influence of emotional factors on learning in physics instruction, International Journal of Science Education 25, 489 (2003).
- Faridi et al. [2021] H. Faridi, N. Tuli, A. Mantri, G. Singh, and S. Gargrish, A framework utilizing augmented reality to improve critical thinking ability and learning gain of the students in physics, Computer Applications in Engineering Education 29, 258 (2021).
- Deslauriers et al. [2019] L. Deslauriers, L. S. McCarty, K. Miller, K. Callaghan, and G. Kestin, Measuring actual learning versus feeling of learning in response to being actively engaged in the classroom, Proceedings of the National Academy of Sciences 116, 19251 (2019).
- Odden and Russ [2019a] T. O. B. Odden and R. S. Russ, Defining sensemaking: Bringing clarity to a fragmented theoretical construct, Science Education 103, 187 (2019a).
- Jennifer Radoff and Hammer [2019] L. Z. J. Jennifer Radoff and D. Hammer, “it’s scary but it’s also exciting”: Evidence of meta-affective learning in science, Cognition and Instruction 37, 73 (2019).
- Odden and Russ [2019b] T. O. Odden and R. Russ, Vexing questions that sustain sensemaking, International Journal of Science Education 41, 1 (2019b).
- Zhao and Schuchardt [2021] F. Zhao and A. Schuchardt, Development of the sci-math sensemaking framework: categorizing sensemaking of mathematical equations in science, International Journal of STEM Education 8 (2021).
- Phillips et al. [2018] A. M. Phillips, J. Watkins, and D. Hammer, Beyond “asking questions”: Problematizing as a disciplinary activity., Journal of Research in Science Teaching 55, 982–998 (2018).
- Sirnoorkar et al. [2023] A. Sirnoorkar, J. T. Laverty, and P. D. O. Bergeron, Sensemaking and scientific modeling: Intertwined processes analyzed in the context of physics problem solving (2023).
- Odden [2021] T. O. B. Odden, How conceptual blends support sensemaking: A case study from introductory physics, Science Education 105, 989 (2021).
- Kapon [2017] S. Kapon, Unpacking sensemaking, Science Education 101, 165 (2017).
- Danielak et al. [2014] B. A. Danielak, A. Gupta, and A. Elby, Marginalized identities of sense-makers: Reframing engineering student retention, Journal of Engineering Education 103, 8 (2014).
- Hunter et al. [2021] K. H. Hunter, J.-M. G. Rodriguez, and N. M. Becker, Making sense of sensemaking: using the sensemaking epistemic game to investigate student discourse during a collaborative gas law activity, Chem. Educ. Res. Pract. 22, 328 (2021).
- Odden and Russ [2018] T. O. B. Odden and R. S. Russ, Sensemaking epistemic game: A model of student sensemaking processes in introductory physics, Phys. Rev. Phys. Educ. Res. 14, 020122 (2018).
- Ford [2012] M. J. Ford, A dialogic account of sense-making in scientific argumentation and reasoning, Cognition and Instruction 30, 207 (2012).
- Russ et al. [2008] R. S. Russ, R. E. Scherr, D. Hammer, and J. Mikeska, Recognizing mechanistic reasoning in student scientific inquiry: A framework for discourse analysis developed from philosophy of science, Science Education 92, 499 (2008).
- Kaldaras and Wieman [2023a] L. Kaldaras and C. Wieman, Instructional model for teaching blended math-science sensemaking in undergraduate science, technology, engineering, and math courses using computer simulations, Phys. Rev. Phys. Educ. Res. 19, 020136 (2023a).
- Kaldaras and Wieman [2023b] L. Kaldaras and C. Wieman, Cognitive framework for blended mathematical sensemaking in science, Int. Journal of STEM Education https://doi.org/10.1186/s40594-023-00409-8 (2023b).
- KUO et al. [2013] E. KUO, M. M. HULL, A. GUPTA, and A. ELBY, How students blend conceptual and formal mathematical reasoning in solving physics problems, Science Education 97, 32 (2013).
- Bain et al. [2018] K. Bain, J.-M. G. Rodriguez, A. Moon, and M. H. Towns, The characterization of cognitive processes involved in chemical kinetics using a blended processing framework, Chem. Educ. Res. Pract. 19, 617 (2018).
- Bing and Redish [2007] T. J. Bing and E. F. Redish, The cognitive blending of mathematics and physics knowledge, AIP Conference Proceedings 883, 26 (2007).
- Kuo et al. [2020] E. Kuo, M. M. Hull, A. Elby, and A. Gupta, Assessing mathematical sensemaking in physics through calculation-concept crossover, Phys. Rev. Phys. Educ. Res. 16, 020109 (2020).
- Gifford and Finkelstein [2020] J. D. Gifford and N. D. Finkelstein, Categorical framework for mathematical sense making in physics, Phys. Rev. Phys. Educ. Res. 16, 020121 (2020).
- Dreyfus et al. [2017] B. W. Dreyfus, A. Elby, A. Gupta, and E. R. Sohr, Mathematical sensemaking in quantum mechanics: An initial peek, Phys. Rev. Phys. Educ. Res. 13, 020141 (2017).
- Sherin [2001] B. L. Sherin, How students understand physics equations, Cognition and Instruction 19, 479 (2001).
- Devlin et al. [2019] J. Devlin, M.-W. Chang, K. Lee, and K. Toutanova, Bert: Pre-training of deep bidirectional transformers for language understanding (2019).
- Note [1] Github repo: https://github.com/tufts-ml/AuSeM.git.
- Thagard and Kroon [2008] P. Thagard and F. Kroon, Hot Thought: Mechanisms and Applications of Emotional Cognition, A Bradford book (MIT Press, 2008).
- Selen and Stelzer [2016] G. G. M. Selen and T. Stelzer, Classical mechanics (FlipItPhysics for University Physics. Macmillan Learning., 2016).
- Cronbach [1951] L. J. Cronbach, Coefficient alpha and the internal structure of tests., Psychometrik 16, 297–334 (1951).
- Amanda Benedict-Chambers and Palincsar [2017] E. A. D. Amanda Benedict-Chambers, Sylvie M. Kademian and A. S. Palincsar, Guiding students towards sensemaking: teacher questions focused on integrating scientific practices with science content, International Journal of Science Education 39, 1977 (2017).
- Watts et al. [2022] F. M. Watts, A. J. Dood, and G. V. Shultz, Developing machine learning models for automated analysis of organic chemistry students’ written descriptions of organic reaction mechanisms, in Student Reasoning in Organic Chemistry (The Royal Society of Chemistry, 2022).
- Long et al. [2024] Y. Long, H. Luo, and Y. Zhang, Evaluating large language models in analysing classroom dialogue., npj Sci. Learn. (2024).
- Ullmann [2019] T. Ullmann, Automated analysis of reflection in writing: Validating machine learning approaches, Int J Artif Intell Educ 29, 217–257 (2019).
- Nakamura et al. [2016] C. M. Nakamura, S. K. Murphy, M. G. Christel, S. M. Stevens, and D. A. Zollman, Automated analysis of short responses in an interactive synthetic tutoring system for introductory physics, Phys. Rev. Phys. Educ. Res. 12, 010122 (2016).
- Wilson et al. [2022] J. Wilson, B. Pollard, J. M. Aiken, M. D. Caballero, and H. J. Lewandowski, Classification of open-ended responses to a research-based assessment using natural language processing, Phys. Rev. Phys. Educ. Res. 18, 010141 (2022).
- Qader et al. [2019] W. Qader, M. M. Ameen, and B. Ahmed, An overview of bag of words;importance, implementation, applications, and challenges (2019) pp. 200–204.
- Murphy [2022] K. P. Murphy, Probabilistic Machine Learning: An introduction (MIT Press, 2022).
- Turner [2024] R. E. Turner, An introduction to transformers (2024).
- Sun et al. [2020] C. Sun, X. Qiu, Y. Xu, and X. Huang, How to fine-tune bert for text classification? (2020).
- Gibaja and Ventura [2014] E. Gibaja and S. Ventura, Multi-label learning: a review of the state of the art and ongoing research, WIREs Data Mining and Knowledge Discovery 4, 411 (2014).
- Bender et al. [2021] E. M. Bender, T. Gebru, A. McMillan-Major, and S. Shmitchell, On the dangers of stochastic parrots: Can language models be too big?, in Proceedings of the 2021 ACM Conference on Fairness, Accountability, and Transparency, FAccT ’21 (Association for Computing Machinery, New York, NY, USA, 2021) p. 610–623.
- Aiken et al. [2021] J. M. Aiken, R. De Bin, H. J. Lewandowski, and M. D. Caballero, Framework for evaluating statistical models in physics education research, Phys. Rev. Phys. Educ. Res. 17, 020104 (2021).
- Tschisgale et al. [2023] P. Tschisgale, P. Wulff, and M. Kubsch, Integrating artificial intelligence-based methods into qualitative research in physics education research: A case for computational grounded theory, Phys. Rev. Phys. Educ. Res. 19, 020123 (2023).
- Fussell et al. [2024] R. K. Fussell, E. M. Stump, and N. G. Holmes, Method to assess the trustworthiness of machine coding at scale, Phys. Rev. Phys. Educ. Res. 20, 010113 (2024).
- Hall and Krist [2025] K. Hall and C. Krist, Unsupervised ml with text data, in Applying Machine Learning in Science Education Research: When, How, and Why?, edited by P. Wulff, M. Kubsch, and C. Krist (Springer Nature Switzerland, Cham, 2025) pp. 281–312.
- Kingma and Ba [2015] D. P. Kingma and J. Ba, Adam: A Method for Stochastic Optimization, in 3rd International Conference on Learning Representations, ICLR 2015, San Diego, CA, USA, May 7-9, 2015, Conference Track Proceedings, edited by Y. Bengio and Y. LeCun (2015).
- Chelba and Acero [2006] C. Chelba and A. Acero, Adaptation of Maximum Entropy Capitalizer: Little Data Can Help a Lot, Computer Speech & Language 20, 382 (2006).
- Xuhong et al. [2018] L. Xuhong, Y. Grandvalet, and F. Davoine, Explicit Inductive Bias for Transfer Learning with Convolutional Networks, in International Conference on Machine Learning (ICML) (2018).
- Masters and Luschi [2018] D. Masters and C. Luschi, Revisiting small batch training for deep neural networks (2018).
- Efron [1992] B. Efron, Bootstrap methods: Another look at the jackknife, in Breakthroughs in Statistics: Methodology and Distribution, edited by S. Kotz and N. L. Johnson (Springer New York, New York, NY, 1992) pp. 569–593.
- Raschka [2020] S. Raschka, Model evaluation, model selection, and algorithm selection in machine learning (2020), arXiv:1811.12808 [cs.LG] .
- Zhao et al. [2024] W. X. Zhao, K. Zhou, J. Li, T. Tang, X. Wang, Y. Hou, Y. Min, B. Zhang, J. Zhang, Z. Dong, Y. Du, C. Yang, Y. Chen, Z. Chen, J. Jiang, R. Ren, Y. Li, X. Tang, Z. Liu, P. Liu, J.-Y. Nie, and J.-R. Wen, A survey of large language models (2024).
- Jiang et al. [2023] A. Q. Jiang, A. Sablayrolles, A. Mensch, C. Bamford, D. S. Chaplot, D. de las Casas, F. Bressand, G. Lengyel, G. Lample, L. Saulnier, L. R. Lavaud, M.-A. Lachaux, P. Stock, T. L. Scao, T. Lavril, T. Wang, T. Lacroix, and W. E. Sayed, Mistral 7b (2023).
- Wojnowicz et al. [2024] M. T. Wojnowicz, K. Gili, P. Rath, E. Miller, J. Miller, C. Hancock, M. O’Donovan, S. Elkin-Frankston, T. T. Brunyé, and M. C. Hughes, Discovering group dynamics in coordinated time series via hierarchical recurrent switching-state models (2024).
- Lerner et al. [2006] R. M. Lerner, W. Damon, E. Thelen, and L. B. Smith, Dynamic systems theories, in Handbook of Child Psychology. Volume 1, Theoretical Models of Human Development (John Wiley & Sons, 2006) 6th ed., pp. 258–312.
Appendix A Physics Problems & Annotations
We provide all four physics problems as they originally appeared to students from Ref. [38]. For each problem, we show the specific annotation scheme - i.e. the specific criteria questions for each problem.
Problem 1
1. Which plot best represents the acceleration curve associated with the displacement and velocity curves shown here?
2. Briefly explain your answer to the previous question.
Annotation Procedure 1
Components
(1) Does the student identify the object (object, car, etc.)?
Characteristics
(2) Does the student identify the orientation (e.g. left or right)?
(3) Does the student identify movement 1 (e.g. slow down)?
(4) Does the student identify movement 2 (e.g. zero)?
(5) Does the student identify movement 3 (e.g. speed up)?
Relationships
(6) Does the student identify the descriptive relationship (e.g position changes is a pos/neg velocity)?
(7) Does the student identify the descriptive relationship (e.g velocity changes is a pos/neg acceleration)?
Problem 2
A physics demo launches a ball horizontally while dropping a second ball vertically at exactly the same time.
1. Which ball hits the ground first?
(a) Dropped ball
(b) Launched (horizontally) ball
(c) Both hit at the same time
2. Briefly explain your answer to the previous question.
Annotation Procedure 2
Components
(1) Does the student identify object 1 (e.g. ball 1)?
(2) Does the student identify object 2 (e.g. ball 2)?
(3) Does the student identify the possible environmental influences on object 1 (e.g. gravity)?
(4) Does the student identify the possible environmental influences on object 2 (e.g. gravity)?
Characteristics
(5) Does the student identify object 1 positioning (e.g. down)?
(6) Does the student identify object 2 positioning (e.g. horizontal up)?
(7) Does the student identify object 1 motion (e.g. speeding up)?
(8) Does the student identify object 2 motion (e.g. speeding up)?
Relationships
(9) Does the student identify the interaction between object 1 and gravity (e.g. gravitational force)?
(10) Does the student identify the interaction between object 2 and gravity (e.g. gravitational force)?
(11) Does the student identify the mechanistic relationship between gravity and a change in motion (e.g. Earth’s gravitational force is a mechanism that causes the object to move down at a constantly increasing speed)?
Problem 3
A girls stands on a moving sidewalk (conveyor belt) that is moving to the right at a speed of 2 m/s relative to the ground. A dog runs on the belt toward the girl at a speed of 8 m/s/ relative to the belt.
1. What is the speed of the dog relative to the girl?
(a) 10 m/s
(b) 6 m/s
(c) 8 m/s
Annotation Procedure 3
Components
(1) Does the student identify object 1 (e.g. girl)?
(2) Does the student identify object 2 (e.g. dog)?
(3) Does the student identify object 3 (e.g. moving sidewalk, conveyor belt)?
Characteristics
(4) Does the student identify the positioning of the dog (e.g. left)?
(5) Does the student identify the positioning of the belt (e.g. right)?
(6) Does the student identify the movement of the girl (e.g. at rest)?
(7) Does the student identify the movement of the dog (e.g. 8m/s constant speed)?
(8) Does the student identify the movement of the belt (e.g. 2 m/s constant speed)?
Relationships
(9) Does the student identify the interaction between the girl and the belt (e.g. moving her to the right)?
(10) Does the student identify the interaction between the dog and the belt (e.g moving the dog to the right)?
(11) Does the student identify the mechanistic relationship between motion of the observer and motion of the observed? (e.g. the speed of the observer relative to the observed object in motion causes the observed motion)
Problem 4
1. The net force on a box is in the positive x direction. Which of the following statements best describes the motion of the box?
(a) Its acceleration is parallel to the x axis.
(b) Its velocity is parallel to the x axis.
(c) Both its velocity and acceleration are parallel to the x axis.
(d) Neither its velocity not its acceleration need to be parallel to the x axis.
Annotation Procedure 4
Components
(1) Does the student identify the object (e.g. object, box, etc.)?
(2) Does the student identify the possible environmental influence acting on the object (e.g. applied net force)?
Characteristics
(3) Does the student identify the object properties (e.g. mass)?
(4) Does the student identify the object positioning before applied force (e.g. upwards)?
(5) Does the student identify the object positioning after applied force (e.g. horizontal)?
(6) Does the student identify the object movement prior to the applied force (e.g. constant velocity)?
(7) Does the student identify the object movement after the applied force (e.g. final velocity)?
Relationships
(8) Does the student identify the interaction between the object and the net applied force (e.g. change in direction)?
(9) Does the student identify the mechanistic relationship between an applied net force and a change in motion? (e.g. a net force applied in the x direction will be a mechanism that causes a change in motion/acceleration in the x direction)
Appendix B Annotation Tables
We provide an annotation summary. We showcase the number of student explanations per problem, and the number of those explanations being labeled as positive for each specific criteria.
| Problem | Number of Responses |
| 1 | 89 |
| 2 | 98 |
| 3 | 100 |
| 4 | 98 |
| Criteria# | Problem 1 | Problem 2 | Problem 3 | Problem 4 |
| 1 | 9 | 89 | 97 | 66 |
| 2 | 4 | 89 | 96 | 62 |
| 3 | 6 | 49 | 78 | 8 |
| 4 | 1 | 48 | 36 | 21 |
| 5 | 4 | 40 | 9 | 46 |
| 6 | 6 | 40 | 80 | 20 |
| 7 | 63 | 46 | 92 | 44 |
| 8 | 46 | 29 | 41 | |
| 9 | 21 | 54 | 42 | |
| 10 | 20 | 36 | ||
| 11 | 18 | 13 |
Appendix C Loss Function Definition
Given the training dataset of assumed independent student embeddings with human annotated labels, we train all parameters to minimize the following loss function:
Here, is the total number of problems, is the total number of student explanations per problem in the dataset , is the total number of domains, and is the set of all context criteria (indexed by ) in a particular domain for a specific problem . In the binary cross-entropy expression, is the ground truth prediction for the student explanation associated with problem with respect to domain and specific criteria . is the corresponding predictive probability.
The loss representing fit to the training data is averaged over all examples. The overall loss also includes a regularization term over all weights. When we train BERTfinetuned, we have both a regularization term in the loss for classifier weights and the encoder weights of BERT . We use an L2SP loss [57, 58], which encourages weights for the LR classifiers to be close to zero, while encouraging BERT encoder parameters to be near pretrained values.
Appendix D Hyper-parameter Strategy
We outline our hyper-parameter procedure step by step such that the experiments in this work can be reproduced. For all experiments, we set the following seeds to control for randomness: when splitting the data, when running experiments, and when bootstrapping test sets.
BoW & BERTfrozen. We randomly select students for our held-out test set ( explanations) and use the rest of the data for a randomized 5-fold cross-validation procedure, where we stratify by student ID. The student ID stratification is to ensure that there is no data leakage between students at train and test time - i.e. we don’t want the model learning student-specific correlations. As such, we ensure that all responses from a specific student are in the same fold. All splits are located in our Github Repo “AuSeM”.
We perform a grid hyper-parameter search while optimizing for the AUROC score over the following: learning rates , and weight decays . We average the value of AUROC from the last epoch from all 5 folds and pick the hyperparameter combination with the best AUROC score. The best model for BoW achieved an AUROC score of with the hyper-parameters: . Similarly, for BERTfrozen, the best model achieved an AUROC score of with the hyper-parameters: .
BERTfinetuned. We use the same held-out test set, but instead of conducting a cross-validation procedure which is computationally expensive for fine tuning, we use the rest of the data as a train/val set for the hyper-parameter selection. The train/val set are also stratified by student ID. The split is located in our Github Repo. “AuSeM”.
We run a hyper-parameter grid search with an Adam optimizer for epochs and batch-size to over the following: learning rate , classifier weight decay , and BERT encoder weight decay . We use two weight decay variables such that we can regularize the weights from BERT separately from the weights of our LR, as we want to best preserve the information learned in the pre-trained model [57, 58]. This totals to 324 fine-tuning experiments for BERTfinetuned on Tufts’ HPC services.
We sort all models by optimal final validation AUROC and select the model that achieves the highest final AUROC . We observe that the model’s behavior appears stable. This model’s hyperparameters are: , , and . Note that prior to sorting by final AUROC, we sorted by AUROC to determine whether a reasonable early stopping point could be identified. While some models did have higher intermediate validation AUROC values, this was only the case for unstable models with too-high learning rates. As such, we determined them to be unreliable, and did not find a place where early stopping would be useful.
Appendix E AUPRC Scores
We show the AUPRC scores for each independent classifier in Fig. 5 and the average AUPRC scores across models in Table 5.
| AUPRC (95% CI) | Runtime | |
| BoW | 0.831 (0.788, 0.870) | 1 min |
| BERTfrozen | 0.881 (0.846, 0.911) | 7 min |
| BERTfinetuned | 0.896 (0.863, 0.927) | 160 min |
Appendix F Confusion Matrices
We present confusion matrices for all three methods. To produce binary predictions, we threshold the predicted probability at a value of 0.5. We observe similar performance here as in the AUROC scores (which take all possible thresholds into account). BERTfinetune has the fewest classification errors overall. However, we see that BERTfrozen makes slightly fewer false positive mistakes (45 compared to 47). For cases where one type of error has a larger cost than another, careful model and threshold selection are important.

### v3

@@
-Using machine learning to measure evidence of students’ sensemaking in physics courses
+Combining physics education and machine learning research to measure evidence of students’ mechanistic sensemaking
 Abstract
-In the education system, problem-solving correctness is often inappropriately conflated with student learning.
-Advances in both Physics Education Research (PER) and Machine Learning (ML) provide the initial tools to develop a more meaningful and efficient measurement scheme for whether physics students are engaging in sensemaking: a learning process of figuring out the “how” and “why” for a particular phenomena.
-In this work, we contribute such a measurement scheme, which quantifies the evidence of students’ physical sensemaking given their written explanations for their solutions to physics problems.
-We outline how the proposed human annotation scheme can be automated into a deployable ML model using language encoders and shared probabilistic classifiers.
-The procedure is scalable for a large number of problems and students.
-We implement three unique language encoders with logistic regression, and provide a deployability analysis on real student explanations from the 2023 Introduction to Physics course at Tufts University.
-Furthermore, we compute sensemaking scores for all students, and analyze these measurements alongside their corresponding problem-solving accuracies.
-We find no linear relationship between these two variables, supporting the hypothesis that one is not a reliable proxy for the other.
-We discuss how sensemaking scores can be used alongside problem-solving accuracies to provide a more nuanced snapshot of student performance in physics class.
+Advances in machine learning (ML) offer new possibilities for science education research.
+We report on early progress in the design of an ML-based tool to analyze students’ mechanistic sensemaking, working from a coding scheme that is aligned with previous work in physics education research (PER) and amenable to recently developed ML classification strategies using language encoders.
+We describe pilot tests of the tool, in three versions with different language encoders, to analyze sensemaking evident in college students’ written responses to brief conceptual questions.
+The results show, first, that the tool’s measurements of sensemaking can achieve useful agreement with a human coder, and, second, that encoder design choices entail a tradeoff between accuracy and computational expense.
+We discuss the promise and limitations of this approach, providing examples as to how this measurement scheme may serve PER in the future.
+We conclude with reflections on the use of ML to support PER research, with cautious optimism for strategies of co-design between PER and ML.
 I Introduction
-Teaching and instruction in undergraduate physics courses has largely relied on problem-solving as the standard method to measure student performance [1, 2, 3, 4, 5, 6].
-Common practice is for “real-time” performance to be measured via multiple-choice or single-solution problems, where canonically correct answers determine the student’s knowledge of the core material.
-Accuracy scores across assignments and examinations, typically coupled with letter grades, act as signals of progress throughout the course as well as final verdicts of student success.
-While engaging in problem-solving is a useful experience for students in a physics classroom, using the problem solution as a measure of student learning assumes a direct correlation that may not always hold.
-Problem-solving accuracy as a measurand assumes that students will engage in a learning process involving the core material to obtain the problem solution.
-Often times, there are alternative strategies for obtaining a problem solution such as rote-memorization of the rules or procedures required for solving similar problem types [7].
-In this scenario, students would score very high on exams that contain these problem types; however given a previously unseen problem structure where the same core material is to be applied, the students would struggle.
-Here, a risk of using problem-solving accuracy as the predominant metric is an inflated sense of confidence in both the instructor and the student that the core material has been learned.
-It could also pose a risk for confounding variables in research studies that aim to investigate how instructional techniques influence student learning [8, 9, 10, 11, 12].
-Over the last fifteen years, researchers in STEM Education (Ed.) have focused more closely on modeling a student’s learning process directly - a process typically referred to as sensemaking [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25].
-Different from memorizing facts, procedures, or problem types, sensemaking is a macro-cognitive process that allows one to construct a model of “how” and “why” with respect to the core material that is consistent with prior experiences and integrates new ones [13].
-In physics courses, there is an emphasis on physical sensemaking, where the goal is for students to construct models that describe physical phenomena.
-This is separate from (blended) mathematical sensemaking, which focuses on how students understand equations and symbolic forms as physical concepts across various scientific disciplines [26, 27, 28, 29, 30, 31, 32, 33, 34].
-Prior work has primarily focused on characterizing physical sensemaking through numerous qualitative analyses of in-moment social dynamics and individual student interviews [25, 23, 14, 15, 16, 17, 18, 19, 20].
-In these scenarios, students were given a problem to solve, but the research objective was focused on modeling the sensemaking process that took place as a problem-solving strategy.
-These empirical investigations resulted in recognizable qualitative outcomes of the physical sensemaking process.
-It is thus a natural next step to consider whether these outcomes can be turned into a post-hoc measurement scheme that reliably quantifies evidence of physical sensemaking given student explanations.
-Crucially, we need such a scheme that scales well with student population size.
-In this paper, we take this next step forward and answer three questions: (1) How can we measure evidence of students’ post-hoc physical sensemaking across multiple problems?
-(2) How can we use machine learning tools to automate this measurement procedure?
-(3) How can these measurements be used to obtain a more informed picture of student performance in a physics course?
-Our main theoretical contribution is a measurement scheme, which quantifies the evidence of students’ physical sensemaking post-hoc by annotating their written explanations for their solutions to physics problems.
-We outline how this human annotation scheme can be transformed into a deployable machine learning (ML) model using language encoders and shared probabilistic classifiers, freeing up time for a human annotator.
-We implement three ML model variations with different text encoders, and provide a deployability analysis on real student explanations to four problems from the 2023 Introduction to Physics course at Tufts University.
-A pipeline based on the BERT language encoder [35] that was fine-tuned for our task achieves the best Area Under the Receiving Operator Characteristic (AUROC) score of on the test set, which contains explanations from students not seen in training data on the same problems as in the training data.
-Lastly, we compute sensemaking scores for all students, and analyze these measurements alongside their corresponding problem-solving accuracies.
-The computed Point-Biserial correlation coefficients between the two variables signify that we cannot claim there exists a linear relationship between problem-solving accuracy and evidence of physical sensemaking in the classroom.
-This demonstrates a need to measure evidence of the sensemaking process explicitly, rather than relying on problem-solving accuracy as a proxy.
-Alongside this manuscript, we contribute practical materials that can be used to help other individuals measure physical sensemaking: (1) examples of implementing the measurement scheme on four physics problems with the corresponding human annotated dataset (385 total student explanations) and (2) ready-to-implement code that generates physical sensemaking measurements for new student explanations on the same set of four problems.
-Both of these materials can be found in the Github Repository “AuSeM” 111Github repo: https://github.com/tufts-ml/AuSeM.git.
-II How to Measure Post-Hoc Physical Sensemaking
-II.1 Defining Physical Sensemaking
-Given a phenomena of interest, physical sensemaking is the macro-cognitive process that constructs the mental model of “how” and “why” for a given physical phenomena by integrating prior and new experiences until consistency is achieved [13].
-Inspired by prior works across research fields that have attempted to theoretically model and empirically characterize student sensemaking [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 37] we summarize the key features of this process: disassembling the phenomena into components, identifying characteristics (i.e.
-feature observations), and considering both descriptive (i.e.
-associations) and mechanistic (i.e.
-cause and effect) relationships.
-During the process, a student may test various hypotheses to “figure things out.” When someone has sensemade, they have formed an organized mental model of these components, characteristics, and relationships that is self-consistent.
-The phenomena can then be explained with natural and mathematical language.
-These prior works, which we discuss in greater detail in Sec.
-II.4, provide the foundational starting point for a quantitative post-hoc measurement scheme.
-Let us consider a problem which contains a physical phenomena to be sensemade as a part of the solution strategy.
-Provided a student’s text explanation of how they solved the problem, a human annotator can ask themselves the following questions:
-1.
-Does the student identify the components that are relevant to the phenomena?
-2.
-Does the student identify the characteristics of the components that are relevant to the phenomena?
-3.
-Does the student identify the descriptive and mechanistic relationships that are relevant the phenomena?
-These initial questions are general and can be further divided into criteria questions that capture more specific outcomes of the sensemaking process.
-Thus, we propose 8 domains that represent what we expect to observe in the student’s explanation as a result of physical sensemaking.
-1.
-Objects.
-The central entities that interact to form the physical phenomena in a single hierarchical level.
-Each object object interaction is relevant to the phenomena.
-Examples: “car”, “box”, “ball”, “people”, “particles”, etc.
-2.
-Influences.
-The non-central environmental entities that interact with objects.
-Only the influence object interaction is relevant to the phenomena.
-Examples: “Earth’s gravitational field”, “human interference”, “electromagnetic field”.
-3. Properties. The relevant qualities that are confined to the physical boundary of an object. Examples: mass, temperature, volume.
-4. Positioning. The spacial states or orientations of an object. Examples: “downwards”, “upwards”, “left”, “right”, “x-direction”, “y-direction”.
-5. Movements. The motion states of an object. Examples: “initial velocity”, “slowing down”, “speeding up”.
-6. Interactions. The exchanges between all object object and influence object pairs. Examples: “gravitational force”, “human pushing”, “electromagnetic force”.
-7.
-Descriptive Relationships.
-The associations between properties, positioning, and movements of an object.
-Examples: “a change in position is a velocity of an object”, “a change in temperature is associated with a change in velocity of a particle”.
-8.
-Mechanistic Relationships.
-The changes that result from object object interactions and influence object interactions.
-Examples: “constant acceleration of a ball caused by Earth’s gravitational field exerting a downward net force”; “acceleration of a book initially at rest caused by a human push”.
-Provided a student’s text explanation as to how they arrived at a particular solution for a problem, the annotator can ask themselves whether the student identifies criteria in each of the 8 domains.
-For example, see Figure 1.
-This is a student explanation in response to Problem 4 [38] below:
+Advances in machine learning (ML) and natural language processing [1, 2, 3, 4, 5] have generated a rush to develop applications in science education, often focused on supporting the analysis and assessment of student thinking [6].
+Much of the promise is to make coding or grading of student work more efficient, as manual assessment of student thinking requires time-intensive, subjective human judgment.
+A key challenge is to provide successful ML support without burdensome requirements for expensive computational or data resources.
+At the same time, scholars argue for patience and care: AI does not provide a ”magic bullet” [7] for research or instruction.
+The benefits of these new technologies will depend on careful investigation, including collaboration and co-design between ML and physics education researchers.
+Examples include recent work by Fussell and her colleagues to propose methodology for assessing the trustworthiness of machine coding, illustrated with analyses of students’ open responses to questions about empirical science, [8].
+In subsequent work, they compare results across different large language models (LLMs) in analyzing skills evident in student lab notes [9].
+We are focusing on the challenge of measuring the students’ sensemaking, which researchers have identified as an essential target for research and instruction, but have been constrained to study in specific episodes or small sets of data [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21].
+These studies took place in a variety of situations–different subject matter, levels of education, forms of data–and there is substantial variation across them in meanings of and nature of evidence for sensemaking.
+Odden and Russ [10] described the literature as “theoretically fragmented.” To resolve this issue, they proposed a definition of sensemaking as “a dynamic process of building or revising an explanation in order to ‘figure something out’ – to ascertain the mechanism underlying a phenomenon in order to resolve a gap or inconsistency in one’s understanding.” We take this as a starting point for our work, recognizing several features we intend ultimately to study: (1) the perception of some gap or inconsistency in one’s understanding; (2) reasoning with the purpose of addressing that gap; and, for science broadly and physics in particular, (3) an expectation of a mechanism underlying the phenomenon [22].
+For now, we focus on developing ML to analyze student reasoning with respect to mechanism, drawing on an earlier analytic framework by Russ and her colleagues for human analysis of mechanistic reasoning [22].
+We review that framework below and discuss the need for its adaptation.
+Designing an approach to ML involves a myriad of contingencies and decisions that depend on the particularities of the situation.
+Here, the particular data we study consist of college students’ brief 1-5 sentence explanations for answers to multiple choice conceptual questions in introductory mechanics.
+The data comes from from a 100-student calculus-based course at Tufts University, which used FlipIt Physics [23] and emphasized students’ learning to engage productively with confusion [24].
+Given this data, two specific questions motivate our current efforts.
+RQ1: How can we refine an analytic framework from physics education research (PER) into a measurement scheme amenable to ML with this data set?
+RQ2: For ML implementations of this scheme, how do design decisions that limit the computational resources needed impact how well the ML-provided labels agree with human coding?
+The remainder of this article is organized as follows.
+Sec.
+II provides background with respect to the data we hope to analyze, prior work on mechanistic reasoning, and prior work to incorporate ML in science education research.
+Sec.
+III presents an account of adapting the framework from Russ et al.
+[22] into a measurement scheme for our particular data set, addressing RQ1.
+Note that we will use the term scheme and the phrase measurement scheme throughout the paper to refer to this adaptation, which is amenable to ML.
+We present the scheme with examples of its application, and we describe the ML in technical detail.
+Part of our purpose in this account is to provide an example of convergence between PER and ML in the co-design of analytic tools.
+Sec.
+IV delves into a specific aspect of the design, RQ2, to compare ML performance across three possible language encoders.
+In this section, we consider both the computational resources required for the model and the level of agreement with a human coder.
+We describe how this approach, which currently depends on problem-specific training, may eventually generalize to support analysis of student responses to new problems without that step of training.
+In Sec.
+V, we discuss the promises and limitations of this use of ML, specifically touching on utility, efficiency, generalizability, reliability, sustainability, and privacy.
+Finally, in Sec.
+VI, we reflect on our efforts to date, considering both our specific aims in developing this particular tool as well as our and the broader community’s more general aims of developing artificial intelligence for applications in education.
+As will be clear, we are not presenting settled science.
+Rather, we are giving examples to illustrate challenges and possibilities of this work.
+To enable others’ explorations, we provide open-access data and code resources in a public Github repository linked at the bottom of page one, under the project name “AuSeM”.
+There, we release our dataset of 385 total student responses across 4 physics problems, each manually-annotated using our approach.
+We also provide a codebook of instructions for human annotators, which reflects updated clarifications that emerged from our inter-rater agreement process.
+These resources could be used to train other models or as a guide for further manual annotation efforts of other data.
+Our provided Python code can be used to reproduce our experiments (re-train each architecture on our data) or generate mechanistic sensemaking scores for new student responses on the same set of four problems.
+II Background
+II.1 The challenge to analyze sensemaking
+The context for this effort is a calculus-based introductory course in Newtonian Mechanics at Tufts University.
+The course places unusual emphasis on students’ sensemaking, with explicit goals of their learning to recognize and to engage proactively with confusion and uncertainty [25].
+The data we examine come from spring 2022, when the course was using FlipIt Physics [23].
+Before each class all semester long, students watched a 20-minute “prelecture” video introducing new concepts, in preparation for the in-person lecture class.
+After the video, students were required to answer a multiple-choice conceptual question to check their understanding.
+For each problem, they submitted an answer as well as a written text explanation of their choice.
+As a motivating example, consider the following problem from the fourth unit of the curriculum:
 Example: Problem 4 1.
@@
 Which of the following statements best describes the motion of the box?
-(a) Its acceleration is parallel to the x axis.
-(b) Its velocity is parallel to the x axis.
+(a) Its velocity is parallel to the x axis.
+(b) Its acceleration is parallel to the x axis.
 (c) Both its velocity and acceleration are parallel to the x axis.
@@
 Briefly explain your answer to the previous question.
-The physical phenomena for a student to sensemake in this problem is an object’s motion under a net external influence.
-Note that our measurement scheme is only made possible by the student’s response to part 2, while part 1 only captures problem-solving correctness.
-We interpret the response as a student’s expression of their problem-solving strategy, and we use it to measure evidence of their physical sensemaking (as shown in Fig 1).
-II.2 Post-Hoc Measurement as Multi-Label Binary Annotation
-Figure 1 demonstrates how an individual student explanation can be annotated with respect to 8 domains; however note that not all 8 domains are relevant (in Fig 1 the descriptive relationships domain is not relevant to the particular problem) and some domains occur multiple times (e.g.
-the object has more than one positioning in the context of the overall problem).
-As such, the full set of yes/no criteria questions across the 8 domains is problem-specific.
-Thus, for a set of problems, an annotator can follow the proposed steps:
-Step 1. Select problems from a particular subject area (e.g. classical mechanics), where there exists a physical phenomena to be sensemade in each problem as a part of the strategy to solution.
+Our challenge is to analyze written text data from student explanations as evidence of their engagement in mechanistic sensemaking. Here are some examples of students’ explanations:
+E1: “According to Newton’s second Law, F(net) = ma, we could know that the direction of a is the direction of F(net)” (Answer b)
+E2: “If the force is acting on it in the positive x direction, then the box must be moving in the positive x direction and will therefore have a parallel acceleration to the x axis since it is not speeding up.
+The velocity would also be parallel because its direction is positive.
+I disagree with myself as I type this but I’m not sure what my opposing argument is.” (Answer c)
+E3: “If a 3 people were pushing the box but one person was pushing harder than the rest, the box would accelerate in the same direction as the greater force.
+However, with velocity, the final velocity has to take into account the initial velocity.
+Though the acceleration might go to the right, if the velocity was going upwards, the box might move at a slant.” (Answer b)
+It is a challenge to decide what to take as evidence of sensemaking, or of not-sensemaking.
+The first makes sense to a physicist, but is it evidence of sensemaking for the student?
+The second does not make sense to a physicist (the box accelerates “since it is not speeding up”?), but the student noted experiencing uncertainty; perhaps it is evidence of their trying to identify that uncertainty.
+Example E3 makes sense to a physicist, and it also shows the student constructing a tangible physical situation.
+Ultimately, we hope to make progress in tapping these data for analysis of sensemaking, to study whether and how students in the course make progress in learning how to learn.
+In this work, we start in that direction by analyzing for evidence of mechanistic reasoning.
+Mechanism is consistently described in accounts of sensemaking in physics [19, 21], and there is a well-established framework for analysis [22].
+II.2 Mechanistic reasoning in science
+Russ and her colleagues [22] adapted work by philosophers of science Machamer, Darden, and Craver [26], who had analyzed the concept of mechanism within areas of biochemistry.
+This led to an analytic framework made up primarily of seven levels organized in a hierarchy:
+#1. Describing the target phenomenon that is observable and reproducible.
+#2. Identifying setup conditions in the spatial and temporal organization of the entities that lead to the phenomenon.
+#3. Identifying entities involved in producing the phenomenon.
+#4. Identifying activities, the “actions and interactions that occur among the entities.”
+#5. Identifying properties of entities that are relevant for producing the phenomenon.
+#6. Identifying the spatial organization of entities that are specifically relevant for the actions and interactions among the entities.
+#7. Chaining backward or forward, in inferences about the entities and activities that could have led to (backward) or could follow from (forward) a given situation.
+These seven categories are logically hierarchical in that higher levels generally require lower levels.
+For example, chaining forward or backward would generally require identifying entities, properties, and activities.
+Russ et al.
+[22] added two categories they considered non-hierarchical:
+Analogies as explicit comparisons to other phenomena.
+Animated models in students’ use of physical actions, whether manipulating objects available or gestures, to simulate the entities and activities they are describing.
+In the 20 years since its introduction, this framework has been applied broadly in PER and other areas of science and engineering education research [27, 28, 21].
+The first example explanation from our data, labeled E1 above, does not include any entities or activities; there is no evidence of mechanistic reasoning.
+That does not mean there was no mechanistic reasoning; only that none is evident in the provided text.
+Nor does it mean the reasoning is bad; it is correct.
+The second explanation E2 gives clear evidence of entities (the box) and activities (its movement).
+Perhaps the student is also thinking of the force as an entity?
+Perhaps, too, there is a chaining, a force causing motion, but the logic is difficult to follow–including evidently for the student.
+It is “moderate evidence of mechanistic reasoning,” by Russ et al’s framework.
+Example E3 has evidence of entities (people, the box), activities (pushing, motion), and of chaining (the box would accelerate in the direction of the greater force/push).
+There is also evidence, although it is not quite as clear, of spatial organization (if the force is sideways, by the people pushing, but the box is moving upward, the box might move at a slant).
+Overall, this example is “strong evidence of mechanistic reasoning.”
+Note that applying Russ et al.’s 7-level general framework to this problem involves both specification and simplification.
+In our data, student responses do not describe a target phenomenon, as the problem prompt does that already.
+Within the identification steps, there are only a small set of possibilities for what could arise as entities and activities.
+As we describe below, in our work to date we take this step, seeking a measurement scheme that can be applied both manually and via ML.
+II.3 Incorporating ML into science education research
+ML tools are now rapidly integrating into PER, predominantly via text classification methods - algorithms that learn patterns in student text responses and sort each response into distinct categories or labels [29, 30, 31, 32, 33, 34].
+This rapid integration has been largely driven by recent advances in large language models (LLMs) [1, 2, 35] trained on vast amounts of language corpora (books, Wikipedia, websites, etc).
+The availability of chatbots like ChatGPT powered by LLMs makes it possible for PER researchers to directly feed text into a chatbot window to be classified [36].
+Below, we contrast different strategies for using LLMs in PER.
+One paradigm for LLM-enabled PER is in-context learning (ICL), where an existing chatbot is provided a student text response and prompted via natural language instructions to classify it into a desired labeling system.
+To be clear, the “learning” here is euphemistic; ICL does not entail retraining the LLM itself at all.
+Instead, it relies on the pretrained ability of the LLM to transform an instructional prompt and provided input text into useful output text.
+ICL is data-efficient, with proponents often providing only a few example demonstrations in the instructions to clarify the desired labeling transformation.
+However, ICL has a number of important limitations.
+First, ICL’s utility relies on the assumption that the LLM is useful for a specific downstream language task in education (correctness, reasoning, sensemaking, etc.).
+The underlying LLM was not trained to do well on these tasks, and it is important to verify at scale whether an LLM is reliable at certain tasks.
+ML researchers have sometimes questioned the utility of ICL, showing that “demonstrations have a marginal impact” [37] and “randomly replacing labels in the demonstrations barely hurts performance” [38].
+Second, ICL is sensitive and variable.
+Outputs can vary depending on the exact prompt [39] and the selection and order of examples [40].
+Even the same exact input may lead to different outputs due to the stochastic nature of chatbots.
+Third, LLMs reflect biases inherent in the source data [41].
+Finally, there is a lack of transparency: closed-source versions such as ChatGPT do not allow users to inspect, understand, or modify how outputs are obtained.
+For these reason, in-context learning methods deserve and require further investigation for their reliability in science education research.
+Other works focus on training supervised ML predictors to accomplish their specific task, integrating accepted human annotation frameworks in science education with machine learning practices [31, 32, 28].
+Watts et al.
+[28] is a prime example of this type of integration between science education and machine learning research.
+Watts et al.
+[27] had previously adapted Russ et al.
+[22] for use in organic chemistry; follow-up work [28] adapts further to apply ML binary classifiers.
+This entailed separating Russ’s original categories into stand-alone binary (present or not present) codes specific to organic chemistry problems.
+For example, they defined the specific code electron movement within Russ’s identifying activities.
+The authors treat each code as independent evidence of mechanistic reasoning, without the hierarchal organization feature of Russ’s framework, which simplifies the ML methodology, a useful first step towards integrating ML and science education research.
+Student responses are used as training data for separate binary classifiers for each code, where various classifier types are attempted, and only the final results of the highest performing classifier are reported.
+Similar to Watts et al, we take the step of adapting the framework to specific features of the tasks in our data set, and we simplify to binary codes, as we describe in the following section.
+We will discuss similarities and differences in technical detail in Sec.
+III.3.
+Fussell et al.
+[8] provide an evaluation of different large-language models for measuring specific skills in students’ typed lab notes.
+The authors compare off-the-shelf pre-trained open-source LLMs with ones that are fine-tuned (further trained) on training data specific to their task.
+The results showed that higher-resource models (e.g.
+more training data, computational power) often outperform lower-resource models for their particular task, although depending on the evaluation metric there are exceptions.
+More specifically, the authors found that all fine-tuned models outperformed the off-the-shelf pre-trained models, and they suggest that for theory-driven classification tasks (i.e.
+where there is a specific task in science education), fine-tuning might be more useful.
+There are a few key distinctions between our current work and Fussell et al.
+Similarly to Fussell et al., we compare the performances of ML model variants in the context of our specific task; however, we do not attempt to provide more general insights about the models on other tasks or datasets.
+We will discuss the similarities and differences between Fussell et al.
+and our work in more technical detail in in Sec.
+III.3.
+III RQ1: PER-ML Measurement Scheme
+III.1 Refining the framework for a particular data set
+Russ et al.
+[22] designed their framework to be applicable across areas and levels of scientific reasoning.
+That breadth of applicability has been helpful for the research community, but its generality presents challenges for the development of an ML tool with context-specific datasets.
+Given a particular data set, such as the conceptual question above about a net force acting on a box, it is possible and advantageous to tune it to the specifics of the task.
+The strengths and limitations of machine learning motivate other simplifications as well.
+Task-specific considerations.
+We are interested in measuring student responses to how they solved a particular problem within a set of closed-ended introductory physics problems.
+Each problem contains its own specific entities, properties, activities, etc.
+that are relevant.
+The box question, for example, mentions only one entity, and there is a limited set of entities (such as people to push) that students might add to the situation.
+As such, the mechanistic evidence that we are looking for is problem-dependent, which allows us to simplify the scheme in several ways.
+(1) No need to re-describe the target phenomenon: Each problem’s instruction text makes the target phenomenon clear; students focused on brief explanations and thus did not restate that phenomenon in their own words.
+For these reasons, we do not include this first code from Russ’s framework in our scheme.
+(2) Restricted set of entities, properties, and activities: The “net force on the box” example problem names only the box as an entity.
+There are in principle other entities one could mention as involved, such as the Earth or the surface under the box, but no students in this data set mention them.
+For each problem in our data, we confine our analysis to a limited, analyst-defined set of entities, properties, and activities that could reasonably be relevant or that arise in the data.
+ML design considerations.
+While we want custom annotations for each problem, developing a separate classifier for each problem would be prohibitively expensive and unlikely to work well without many training examples.
+We are thus interested in combining powerful language encoders [1, 3, 35, 2] with shared classification heads that take advantage of similar word occurrence and grammatical structure across problems to efficiently annotate new examples.
+As such, it is important to construct multi-label classification categories that share similar linguistic features across student responses and problems.
+This influences three key design decisions.
+(1) Share initial and subsequent spatial organizations: We assume a description of the spatial organization of entities would contain similar semantic and structural features independent of whether it refers to the initial, intermediate, or final conditions.
+These temporal states can be represented by a separate context embedding (further discussed in Sec.
+III.3), such that evidence for each temporal state can still be individually annotated, but spatial linguistic features can be shared.
+(2) Separate central from non-central entities: By definition, central entities are connected to the other codes (properties, activities, etc.).
+As such, central and non-central entities will contain dissimilar relationships among words in a student response.
+Keeping these categories separate is especially useful for future work that incorporates dependencies between domain labels.
+(3) Separate movement activities from interactions: Descriptions of individual entity movement activities (e.g.
+the object is moving with an initial velocity) are different from descriptions of the interactions between entities (e.g.
+the hand pushes the book) both semantically and structurally within a sentence.
+Like Watts et al.
+[28], we relax the hierarchy of Russ’s framework, coding only for the presence or absence of evidence in each category.
+This simplifies our annotation scheme and ML method for now; later in Sec.
+VI we discuss plans going forward to reintroduce some form of that hierarchy.
+This would be an improvement on our current modeling assumptions: for example, evidence of properties is likely to correlate with evidence of entities.
+Proposed domains. Given the above considerations, we propose 7 domains, broad categories of yes/no questions that measure evidence of mechanistic sensemaking in student thinking via the framework:
+1.
+Objects.
+Identifying the central entities that interact to comprise the physical phenomena in a single hierarchical level.
+These objects have properties, spatial layout, and movement relevant to the solution process.
+Examples: “car”, “box”, “ball”, “particles”, etc.
+2.
+Influences.
+Identifying the non-central environmental entities that affect the objects and their activities.
+Only the influence object interaction is relevant to the phenomena.
+Examples: “Earth’s gravitational field”, “human hand”, “electromagnetic field”.
+3.
+Properties. Identifying the relevant qualities that are confined to the physical boundary of an object. Examples: mass, temperature, volume.
+4.
+Spatial organization. Identifying the spatial states or orientations of an object. Examples: “downwards”, “upwards”, “left”, “right”, “x-direction”, “y-direction”.
+5.
+Movement activities. Identifying the motion states of an object. Examples: “initial velocity”, “slowing”, “speeding”, “accelerating”.
+6.
+Interactions. Identifying the physical exchanges between all object object and influence object pairs. Examples: “gravitational force”, “human pushing”, “electromagnetic force”.
+7.
+Chaining.
+Identifying the characteristic (a property, spatial organization, or movement activity) changes that result from specific object object interactions and influence object interactions.
+Examples: “constant acceleration of a ball caused by Earth’s gravitational field exerting a downward net force”; “acceleration of a book initially at rest caused by a human push”.
+Our scheme corresponds with Russ’s framework as follows: the properties domain directly maps to Russ’s category #5; the objects domain and influences domain combine to obtain #3; the spatial organization domain encompasses #2 and #6; the movement activities and interactions domains combine to form #4, and the chaining domain maps to #7.
+We do not consider stating the target phenomena, analogies, or mental model animation in this work - although these are interesting additions for future research.
+Example problem and associated coding.
+Consider again the “force on a box” problem from Sec.
+II.1.
+An example student explanation is shown in Fig.
+1, with specific parts color-coded by our 7 domains.
+This coding demonstrates how parts of an individual student response can be associated to each of the 7 domains.
+These specific parts could be later used to answer yes/no questions to quantify evidence of sensemaking, as described in Sec.
+III.2.
+Some domains correspond to multiple parts of the response (e.g.
+the object has more than one spatial organization in the context of the overall problem).
+Note that not all 7 domains need to be coded for if some are irrelevant to a problem, though all domains are relevant to the example problem here.
+III.2 Measurement as multi-label binary annotation
+Given a specific physics problem, our measurement scheme requires the annotator to develop a list of yes/no criteria questions for each of the 7 domains that is problem-specific.
+Thus, given a dataset of potential problems and student responses to be measured, an annotator can follow the proposed steps:
+Step 1. Select problems from an introductory physics course on classical mechanics, where there exists a physical target phenomena to be understood.
 Step 2.
 For each problem and for each domain , determine the set of yes/no criteria questions that are pertinent to the specific problem.
-For example, in Problem 4 (II.1), there is only a single object and thus with the criteria question Does the student identify the object?.
-However, there are two positioning criteria questions that are relevant to sensemaking an object’s positioning under a net external influence: Does the student identify the object positioning before the net applied force?
-Does the student identify the object positioning after the net applied force?
-The annotator should end up with the total number of criteria questions per problem as .
+For example, in Problem 4 from Sec.
+II.1, there is only a single object, and thus with the criteria question Does the student identify the object?.
+However, there are two spatial organization criteria questions : Does the student identify the object spatial organization before the net applied force?
+Does the student identify the object spatial organization after the net applied force?
+We denote the total number of criteria questions for problem as .
+The full list of criteria for Problem 4 is displayed in the Appendix A.
 Step 3.
@@
 To further compute a sensemaking score one can sum the number of and divide by the length of the vector .
-This score is to be interpreted as the amount of evidence of physical sensemaking.
-An annotator can also decide to weight the domains (or specific criteria) differently.
-In this work, we consider all domains and criteria to equally contribute to the sensemaking score.
-To finish off our example, the criteria vector for the student explanation in Fig.
-1 is with the corresponding sensemaking score .
-The student achieved such a high score by imagining a scenario where multiple interactions are influencing the boxes’ dynamics and communicating the resulting mechanistic relationship between a net external force and an the boxes’ change in motion.
-The student received a yes for all criteria except for the one in the property domain.
-The full list of criteria for Problem 4 is displayed in the Appendix A).
-II.3 Theoretical Benefits and Limitations
-Utility.
-This annotation scheme is intentionally designed to measure evidence of physical sensemaking given student explanations.
-The measurements allow for evidence of sensemaking to be quantitatively compared to other numerical values of interest, and distinguished from correctness.
-For example, evidence of sensemaking is different from measuring sensemaking correctness (i.e.
-the student is engaging in the sensemaking process with information that is determined by the authority figure to be canonically correct).
-The current scheme could be easily modified for this other purpose; the annotator would simply need to choose to annotate for both sensemaking and correctness according to the problem (e.g.
-the student identifies the correct positioning).
-However in this work, we stick to sensemaking scores that measure evidence of sensemaking, and leave correctness for future studies.
-Moreover, we emphasize that low sensemaking scores measured do not indicate that students cannot sensemake.
-It simply measures that in the current snapshot, evidence of sensemaking is low.
-There could be several reasons as to why the evidence is low (i.e.
-the problem itself didn’t encourage sensemaking, the student applied a memorized fact or procedure, etc.).
-Put concretely, our annotation scheme does not allow one to make any claims about causality.
-Flexibility.
-Individuals can design or select their own problem sets, and then annotate the student explanations according to the same 8 domains.
-In proposing such an adaptive scheme, we attempt to balance imposing a systematic procedure with allowing the annotator to have enough freedom to make decisions.
-Reproducibility.
-We provide a step-by-step outline of the annotation procedure, concise definitions of the 8 domains, and annotated examples of 385 student explanations across 4 physics problems of different criteria vector lengths.
-Our hope is that this content leaves little ambiguity for a new annotator looking to reproduce our results or implement the scheme in their own work.
-However, we acknowledge that some variability among annotators could exist and be quantified in the future with inter-rater reliability studies [39].
-Reducibility.
-Each criteria vector length is the amount of binary information (bits) required to represent the evidence of a student’s sensemaking with respect to a particular problem.
-It is indeed possible that the 8 domains and criteria could be reduced such that a fewer number of binary variables are needed.
-While a reduced model form is not currently known; we speculate that such a form could exist.
-Scalability.
-We acknowledge that manual annotation of a large number of student responses would require many hours of labor, often infeasible in practical settings.
-For example, the student explanation in Fig 1 took minutes to annotate.
-For a class of one-hundred students, this extrapolates to annotation hours with respect to a single problem.
-To resolve this scalability challenge, we show how the procedure can be automated by training an ML classification model on student explanations.
-II.4 Sensemaking Related Works
-Odden and Russ explicitly define sensemaking as “the dynamic process of building an explanation to resolve a perceived gap or conflict in knowledge.” The authors provide examples of individuals engaging in this process, and discuss how sensemaking can be conceptualized in three ways: a stance or framing (i.e.
-an individuals’ expectation of themselves in their current situation), a cognitive process built on the theory of knowledge integration, and a discourse practice related to argumentation.
-Each conception addresses what sensemaking is with respect to a different system level of study.
-Cognitive theories are based on the mental components and interactions happening during sensemaking, whereas framing can be considered one hierarchal level up as a latent description of a student’s current expectations in their environment when sensemaking is taking place.
-Lastly, discourse analysis is based on studying the content within and form of conversational dialogue between students.
-Qualitative analysis has shaped our understanding of how sensemaking begins, functions, and ends for students.
-Odden and Russ [15] demonstrate how vexing questions stimulate the sensemaking process and can sustain it for longer durations of time than typically observed.
-Benedict-Chambers et al.
-[40] provides four specific types of questions focused on guiding students towards sensemaking including explication, explanation, science concept, and science practice questions.
-To characterize the explicit features of in-moment sensemaking, Odden and Russ [23] propose that the process be modeled as an epistemic game.
-This model introduces sensemaking as a three step procedure: 1.
-noticing a gap or inconsistency, 2.
-generating a mechanistic explanation, and 3.
-resolving the gap or inconsistency.
-The authors discuss a precursor to sensemaking, which includes assembling a knowledge framework by initially discussing ideas socially.
-This brainstorming step is necessary for one to eventually recognize a gap in their own knowledge system.
-During the game, the students can make moves (e.g.
-refining a question, appealing to reality, building an analogy, etc.) and if certain constraints are violated, students exit the game play.
-Three insightful constraints are discussed: 1.
-sensemaking is a discussion between individuals rather than an imbalanced monologue, 2.
-sensemaking occurs when there is a phenomena to sensemake that has not already been sensemade, and 3.
-sensemaking contains self-generated definitions of ideas alongside formal definitions.
-Kapon [20] puts forth a metric for how explanations generated during the sensemaking process are implicitly evaluated by individuals - based on intuitive knowledge, mechanism, and framing.
-More specifically, students evaluate explanations as “good enough” based on whether it aligns with their phenomenological experiences, is mechanistic rather than teleological, and satisfies their immediate expectations of their own understanding.
-Odden [19] explores how conceptual blends support sensemaking through a qualitative case study of students using concepts from kinematics to understand the mechanisms that underlie the phenomena in electric circuits.
-The author proposes the approach of generalizing conceptual ideas from one sub-area of physics to another as a mechanism that allows students to make connections in the sensemaking process.
-Furthermore, Sirnookar et al.
-[18] explores the role of modeling in sensemaking and analyzes the interactions of these two processes as two students are interviewed while solving a physics problem.
-The authors find that barriers in modeling can inhibit students’ sustained sensemaking.
-In pursuit of the longer-term impacts of sensemaking, Radoff et al.
-[14] provides a qualitative account of a student engaging in sensemaking throughout an academic semester.
-Interviews and coursework demonstrated that experimentation with sensemaking caused more positive associations with learning, taking on challenges, and constructive confusion.
-In Kuo et al.
-[31], the authors show through a controlled classroom experiment that students who are instructed with mathematical sensemaking stimulation are prone to obtaining solutions to physics problems that are more accurate and insightful than those who did not receive this type of explicit instruction.
-The sensemaking literature is fragmented across several scientific disciplines and is also typically broken into categories of scientific and mathematical sensemaking, where more recent work has attempted to define a blended model of the two [16, 26, 27].
-Kaldaras and Wieman [27] build upon prior works to put forth a cognitive model of blended math-science sensemaking that introduces three levels of sophistication: 1.
-developing qualitative relationships among relevant variables in mathematical equations that describe a particular phenomenon, 2.
-developing mathematical relationships among these variables, and 3.
-explaining how the mathematical operations used in the equation connect back to the physical phenomena.
-The authors attempt to validate their model by analyzing student responses from interviews, specifically probing the various levels of sophistication in the framework.
-While the authors found convincing evidence that the levels of sophistication were present across all student responses, it remained inconclusive as to whether there exists an actual sophistication ranking in the levels.
-As a result, it is still uncertain whether there exists a hierarchy among cognitive connections in math-science sensemaking.
-This work is directly linked with research that asks: what does it mean to understand a physics equation?
-[34] and seeks out structure in symbolic expressions that aid in student understanding.
-The works above collectively agree on a key characteristic of the sensemaking process: the dynamic generation of mechanistic explanations for a given phenomena.
-When students have resolved their inconsistency, they have settled on a mechanistic explanation.
-The emphasis on mechanism in sensemaking is historically tied with research that characterizes student’s mechanistic reasoning in the classroom [17, 25, 41].
-Russ et al.
-is arguably the the most influential contribution made thus far in defining mechanism in the context of STEM Ed., providing a framework for discourse analysis developed from ideas in philosophy of science [25].
-In the framework, mechanistic reasoning is characterized as a process that uses experience-based evidence to construct causal relationships among objects describing “how underlying structure and activities can account for observable changes in a system”.
-The work puts forth a coding scheme of nine categories: 1.
-describing the target phenomenon they are trying to explain, 2.
-identifying setup conditions of the spatial and temporal organization of entities that construct the phenomena, 3.
-identifying the entities as components that construct the phenomena, 4.
-identifying activities (actions and interactions) of the entities, 5.
-identifying properties of entities, 6.
-identifying the spatial organization of entities relative to each other, 7.
-chaining reasoning strategy for backward or forward inference, 8.
-analogies, and 9.
-mental model animation.
-This framework has been widely adopted by the STEM Ed.
-community for qualitatively analyzing student reasoning, and we highlight that it overlaps significantly with our 8 domains presented to measure the evidence of sensemaking.
-There are a few key differences.
-The first is that we separate entities into objects and influences, where influences interact with objects in such a way that only the influence object effect is relevant to the particular phenomena.
-This is especially useful when dealing with fields in physics, which are impacted from object interactions, but the effects are oftentimes ignored in classical settings or introductory courses.
-Our two domains positioning and movement track both the set-up and subsequent conditions of the dynamics of each object.
-The positioning domain also includes how the objects relate to one another structurally (i.e.
-object A is on top of object B) when this is relevant for the particular phenomena.
-As such, we reduce three of the codes (2,6,9) in Russ et al.
-to a single domain.
-As we aim to measure post-hoc explanations rather than in-moment discourse, we can code for the chaining and analogy creation process (which can be viewed as sub-sensemaking processes) proposed in Russ et al.
-by annotating whether the student identifies the relevant mechanistic relationships.
-In addition, we add a domain for descriptive relationships, which are the relations between the properties of objects - more akin to correlation than causal mechanism.
-Otherwise, with respect to Russ et al., our properties domain directly maps to code 5 and our interactions domain directly maps to code 4.
-As previously stated, our work builds off of the numerous qualitative studies discussed above that have identified the features of sensemaking and mechanistic reasoning through analyzing student discourse.
-We view our main contribution as transforming these features into 8 domains that contain sub-questions with binary outcomes such that we can obtain a post-hoc sensemaking measurement of a student’s explanation with respect to any particular physics problem.
-Introducing such a binary encoding scheme allows one to quantify evidence of student sensemaking and naturally lends itself to automating the measurement scheme with ML.
-III How to Automate Measurements
-III.1 Multi-Label Binary Classification Model
-ML models have been shown to be capable of replacing labor-intensive encoding of qualitative data with automated procedures [41, 42, 43, 44, 45].
-In this work, we only consider models which are open-access, which allows us to train classifiers for our specific task and does not require us to share student data with a 3rd party.
-Furthermore, we limit our consideration to models which are capable of running on consumer hardware.
-In our scenario, the task is the same as the human annotator’s: to map student text explanations across independent problems to sensemaking scores .
+This score is to be interpreted as the amount of evidence of mechanistic sensemaking.
+An annotator could decide to weight the domains (or specific criteria) differently.
+In this work, we keep the weights equal.
+To finish our example, the criteria vector for the student response in Fig.
+1 is with the corresponding sensemaking score , a “yes” code for all domains but “property.” Considering the response itself, a high score seems appropriate: The student imagined a tangible scenario with multiple influences on the box.
+III.3 Multi-label binary classification model
+Here, we describe an ML model we constructed for our task, that can automatically map a student text response to a numerical sensemaking score, where a higher value indicates greater evidence of mechanistic sensemaking.
+While our approach relies on problem-specific criteria, some aspects of the model are shared across problems to gain statistical strength.
+The overall model is modular: components like the encoder can be varied to tradeoff between model capacity and resource consumption.
 From text to embeddings.
-Given a student explanation , we can select an encoder to map all text in the response to a numerical representation .
-A simple encoder design is counting; we determine a vocabulary as the set of words across all student explanations in the training dataset, with an additional vocabulary entry for unknown words, and record the number of times each one appears in the example.
-This numerical representation, or embedding , typically called BagofWords, reduces the richness of the text explanation to the pattern of word occurrence [46, 47].
+Given a student response , we can select an encoder to map all text in the response to a numerical representation .
+A simple encoder design is counting; we determine a vocabulary as the set of words across all student responses in the training dataset, with an additional vocabulary entry for unknown words, and record the number of times each one appears in the example.
+This numerical representation, or embedding , typically called Bag-of-Words, reduces the richness of the text explanation to the pattern of word occurrence [42, 43].
 From here on, we will refer to this encoder as BoW.
-To capture more complex patterns (e.g relative co-occurrence between all words in the explanation, in-context word occurrence, and word positional associations with grammar rules), one can use a large language model (LLM) encoder [48].
+To capture more complex patterns (e.g relative co-occurrence between all words in the response, in-context word occurrence, and word positional associations with grammar rules), one can use a large language model (LLM) encoder [44].
 For example, Google’s BERT (Bi-directional Encoder Representations from Transformers) model combines tokenization, self-attention, and a multi-layer perceptron neural network - all pre-trained on English Wikipedia and BookCorpus to generate token embeddings [1].
-A token is a set of characters containing a unique ID in BERT’s vocabulary, such that an entire text can be converted to tokens with individual embeddings , which carry information about all other tokens in the sentence.
+A token is a set of characters containing a unique ID in BERT’s vocabulary, such that an entire text can be converted to tokens with individual embeddings , which carry information about all other tokens in the passage.
 BERT also uses a sentence-level embedding, known as , which can be used as a full-response representation rather than using distinct representations for each token.
 From here on, we will refer to this encoder as BERTfrozen.
-Additionally, can be fine-tuned by training BERT’s previously trained weights on new downstream classification tasks [49].
+As a final encoder option, the sentence embedding can be fine-tuned by training BERT’s previously trained weights on new downstream classification tasks [45].
 We will refer to this encoder as BERTfinetuned.
-While requiring more computational resources, this allows for the embedding to capture patterns in the text that are specific to the target prediction task.
+This allows for the embedding to capture patterns in the text that are specific to the target prediction task.
+It requires significantly more computational resources, however.
+Current LLMs, such as Google’s Gemini or OpenAI’s ChatGPT, require far more still.
+Part of our interest is to minimize the computational demands of the ML.
 Shared classifiers across problems.
-Given a generated embedding for a student explanation, we want to classify the representation into its correct criteria vector .
+Given an embedding representation for a student response, we want to correctly assign it to a criteria vector .
 Thus, we have a multi-label task [46].
-If we construct separate binary classification models for each criteria question, then we are required to train independent classifiers.
-Rather, we propose to work off of an informed assumption: shared domains will have similar word occurrence and grammatical structure across P problems.
-For example, the phrase “speeding up” should indicate evidence of sensemaking in the movement domain regardless of the problem it appears in , and words that should be classified as objects will share similar positions in the sentence across problems due to object-subject-verb rules.
-Thus, for problems, we can train 8 independent classifiers.
+If we were to construct separate binary classification models for each criteria question, then we would be required to train independent classifiers.
+Rather, we propose to work off of an informed assumption: shared domains will have similar word occurrence and grammatical structure across problems.
+For example, the word “speeding” should indicate evidence of sensemaking in the movement activities domain regardless of the problem.
+Furthermore, words that should be classified as objects will share similar positions in the sentence across problems due to object-subject-verb rules.
+Thus, for problems, we propose to train 7 classifier heads, one for each domain.
 We reduce from a scaling linear in to constant in .
-Note that we are proposing to train 8 classifiers to obtain annotations for student explanations across problems, where each problem has a different number of criteria questions per domain .
-Thus, to obtain an annotation for each criteria question within a domain, we add a criteria context embedding to each text embedding for a given student explanation.
-The updated embedding vector is denoted as .
-The index specifies the criteria question that we would like for the classifier to label (e.g.
-Does the student identify the object positioning before applied force (e.g.
+In order for this shared set of classifiers to effectively annotate multiple problems each with multiple yes/no criteria questions, we design each classifier to take an additional input beyond the embedding of student response text: a criteria context embedding that encodes the text of the yes/no question being asked (e.g.
+Does the student identify the object spatial organization before applied force (e.g.
 upwards)?).
-We note that there are multiple possible ways to encode this context.
-For example, rather than adding the vectors to obtain the updated embedding, one could concatenate the two vectors together.
-Additionally, one could embed the entire problem text alongside the specific criteria question.
-As a starting point, we’ve selected our approach of adding the embeddings together and only including the criteria question.
+The index specifies the criteria question the classifier should label.
+Each classifier must then take in student response embedding and criteria context embedding .
+There are multiple possible ways to combine the two inputs into a joint representation of response and context vectors that is the ultimate feature vector for the classifier.
+As a starting point, we have selected a simple approach of adding the embeddings together: .
+Possible alternatives would be to concatenate the two vectors together, or to produce one embedding from the combined text of question and response.
 We leave the exploration of alternative methods for future work.
-Figure 2 visually illustrates how the model makes a probabilistic prediction answering a specific yes/no criteria question given a student-generated text response.
+Fig.
+2 visually illustrates how the model makes a probabilistic prediction for a specific yes/no criterion given a student’s response to a question.
 Given embedding , we use a domain-specific logistic regression (LR) classifier to produce the predicted probability .
 We set , where are the weight vector and bias scalar for domain , and denotes the logistic sigmoid function.
-The figure also shows how that prediction impacts the loss used to train the model.
-For training, we have a dataset student embeddings with human annotated labels (the criteria vectors) .
-Each represents a single binary label with respect to criteria question of domain for the student explanation .
+The figure also shows how that prediction impacts the loss, which measures the difference between the model’s current predictions and human coding.
+For training, we have a dataset of student embeddings with human annotated labels (the criteria vectors) .
+Each represents a single binary label with respect to criteria question of domain for the student response .
 Each is assumed conditionally independent of other labels given and all model parameters.
-We minimize the following binary cross entropy loss with an added regularization term over the weights averaged over all explanations, domains, and criteria:
+We minimize the following binary cross-entropy loss
+This loss is aggregated over all explanations , domains , and criteria .
+We also include an added regularization term that penalizes the sum-of-squares of the weight parameters.
 For the precise loss notation utilized in our methods including regularization terms, see Appendix C.
-The training procedure encourages all 8 classifiers to accurately predict their assigned pieces of the criteria vector for an embedded student explanation .
+The training procedure encourages all 7 classifiers to accurately predict their assigned pieces of the criteria vector for an embedded student response .
 For some encoder choices, we do not update any encoder parameters during training; only the classifier weights and biases are updated via gradient descent.
 When fine-tuning BERT, our gradient descent procedure updates encoder parameters alongside classifier weights.
-Both are explored in our deployability analysis in IV.1.
-III.2 Theoretical Benefits and Limitations
-Utility.
-The proposed model maps new student text explanations across a set of problems to sensemaking scores , freeing up time for an annotator.
-However, the current version of this model is limited to measuring sensemaking scores of student explanations on the same set of problems.
-We hypothesize that with more labeled training data across new problems, the model proposed here will have enough data to perform well on unseen problem instances in physics.
+Both are explored in our model comparative analysis in Sec.
+IV.
+Technical comparison with related work.
+Similar to Fussell et al.
+[8], we train a BagofWords (BoW) encoder and a BERTfinetuned encoder.
+We also observed that the BERTfinetuned encoder achieves better performance than the BoW encoder for our specific task.
+In general, our later results support their conclusion that fine-tuning adds value for a theory-driven classification in PER.
+They also compared two model architectures with multiple LLaMa encoders [2], which require significantly more computational resources than our BERT encoders.
+Similar to Watts et al.
+[28], we train multiple binary classifiers for task-specific evidence of mechanistic reasoning.
+They determined convolutional neural networks (CNNs) to outperform other model architectures such as naive Bayes, logistic regression, support vector machines, and transformers.
+Their results show high accuracy on the test data with slightly lower Cohen’s Kappa values to account for the data imbalance.
+We chose to continue with BERT encoders, expecting better inductive bias for our task than CNNs, especially when pretrained on large language corpora then fine-tuned on our limited data.
+IV RQ2: How ML Design Choices Can Impact Performance
+We seek a performance comparison of the proposed classification models outlined in Sec.
+III.3 with different encoding and training strategies.
+More specifically, our intention is to estimate how these models will perform given explanations from new students that it has never seen before, but on the same set of problems.
+Below, we outline our implementation and evaluation procedures, then present and discuss results.
+Dataset of student responses.
+Our dataset consists of 385 total de-identified student-responses across four problems distributed to 100 students in the first few lectures of the 2022 Tufts University Introduction to Physics course.
+The retrospective analysis of this deidentified human subjects data was approved by our institutional review board.
+All problems have the same structure as Problem 4 in Sec.
+II.1 shown above and are presented to the students as pre-lecture digital assignments [23].
+See Appendix A for exact problem text.
+The same students are asked to solve each problem and explain how they obtained their solution.
+A few students chose not to respond to a particular problem, which brings our total number of explanations to .
+Annotating ground-truth labels.
+Complete annotation for all 4 problems was carried out by a domain expert: the first author of this work who holds a doctorate in physics.
+This included defining the yes/no criteria questions for each problem (provided in Appendix A) and assigning binary values for each student response.
+These problems vary in criteria vector size: .
+Some descriptive statistics summarizing annotations across problems are provided in Appendix B.
+We validated inter-rater reliability post-hoc, as described in Sec.
+V.3.
+In future work, we plan to incorporate training examples from multiple expert annotators.
+Experimental setup.
+To evaluate model performance on never-before-seen responses, we randomly select 15 students who responded to all four questions to use as a held-out test set.
+We use the rest of the data (325 samples) for training and validation.
+To prevent data leakage, we stratify the data by student to ensure that all responses from the same student across 4 problems end up either exclusively in the training set or exclusively in the test set.
+Using this train/test split, we fit the models BoW, BERTfrozen, and BERTfinetuned described earlier.
+All models are trained with the Adam optimizer [47] for epochs.
+Convergence was verified by observing loss function behavior.
+We perform a grid search over key hyperparameters: learning rate and regularization strength ( for classifier weights and for BERT model weights); reproducible details are available in Appendix D.
+For BoW and BERTfrozen, we use 5-fold cross-validation, also stratified by student, to select hyperparameters.
+We pick the combination with the best average AUROC score across folds.
+The primary performance metric we report and evaluate is the AUROC, the Area Under the Receiver Operating Characteristic (ROC) curve.
+For a given classifier, this curve depicts the true positive rates (y-axis) and false positive rates (x-axis) achieved across different numerical thresholds for turning predicted probabilities into binary decisions.
+There are two main reasons we believe AUROC to be the most useful metric for our context: First, AUROC does not require us to choose an arbitrary probability threshold to distinguish label 1 predictions from label 0 predictions.
+Instead AUROC sums over all possible thresholds, providing an overall numerical summary where increasing values indicate better performance.
+AUROC is bounded in .
+Second, due to imbalanced annotations in our training dataset (i.e.
+some multi-label tasks have more of the positive or negative class rather than a 50-50 split), AUROC is an easier to interpret metric to report than precision or accuracy, as its chance baseline is independent of imbalance.
+For these reasons, we use AUROC as the primary performance metric throughout our analyses.
+For the interested reader, we also include additional metrics like area under the precision-recall curve in the Appendix E and confusion matrices in Appendix F.
+Fine-tuning details.
+For BERTfinetuned, we require a larger amount of computational resources per experiment.
+Our grid search over hyperparameters entails 324 separate efforts to train this encoder’s millions of weights.
+We thus used a single training and validation set to run the hyperparameter grid search, avoiding the extra costs of 5-fold cross-validation.
+To better preserve the prior information learned in BERT’s original training, we use L2SP regularization [48, 49] when training BERTfinetuned.
+This means we enforce two separate sum-of-squared-difference penalties in the loss, one for BERT encoder weights and another for the LR classifier head weights.
+Each has a different desired mean and a different strength hyperparameter that is tuned.
+We set the batch-size to , which has been shown empirically to work well with deep neural networks [50].
+During finetuning, we originally incorporated an domain classifier that we now believe to be out of scope for this particular work.
+As the domain-specific classifiers are independent, this does not influence the results of BoW or BERTfrozen.
+However, we acknowledge this may impact our reported performance of BERTfinetuned as the embeddings across classifiers are shared.
+We chose not to retrain as the model resources required to do another hyperparameter optimization (another set of 324 HPC experiments) is high.
+Re-training without the domain would likely only increase the performance of BERTfinetuned on 7 key domains, and as we show below it is already performing better than other variants .
+Results from initial validation phase.
+The best model for BoW achieved an average AUROC score of across validation folds with the hyper-parameters: .
+Similarly, for BERTfrozen, the best model achieved an AUROC score of with the hyper-parameters: .
+For BERTfinetuned, after running 324 fine-tuning experiments on Tufts’ HPC services, we picked a model with final AUROC of on the validation set, which had the following hyperparameters: , , and .
+Results on heldout test data.
+We retrain our selected models (with model-specific optimal hyperparameter settings) on all training and validation data.
+We then evaluate on the held-out test set of 60 responses from randomly selected students.
+In Fig.
+3, we show the average AUROC across independent test predictions for each of the 7 domain classifiers and for each model type.
+To create approximate uncertainty intervals, we use bootstrap resampling of our test set [51, 52] to generate same-sized test sets that represent plausible unseen student responses, and compute the AUROC of each one.
+Intervals in Fig.
+3 depict the 2.5-97.5th percentile range of these bootstrapped AUROCs.
+Beyond the per-domain results in Fig.
+3, we show the overall average AUROC for each model in Table 1.
+This micro-average is computed by weighting each annotation label equally across all domains.
+As such, these values represent the annotation quality of the multi-task classification set-up as a whole for each model.
+In Table 1, we also report the time taken to train the final selected model.
+To assess the significance of differences between models, we provide the confidence interval in Table 1 in the difference in AUROC between models, computed by taking percentiles from our bootstrapping procedure.
+These intervals can be interpreted as indicating uncertainty in AUROC difference across models.
+When comparing model A and B, if the interval is all positive and an exact zero is outside of the interval, we can be confident that there exists an actual positive difference between models, and model A is indeed better than B.
+BoW performs decently overall with an AUROC score of and trains efficiently in 1 minute.
+For the objects domain in particular, it performs nearly as well as the other two models.
+BERTfrozen performs slightly better with an AUROC score of and takes 7 minutes to train.
+It performs the best for two domains: objects and chaining.
+It is able to reliably outperform BoW on average, as zero in not within the confidence interval for AUROC difference reported in Table 1.
+(The interval containing zero would indicate that no significant difference between models is plausible.) BERTfinetuned achieves an AUROC score of .
+It performs the best for influences, spatial organization, movement activities, and interactions.
+It takes significant time to train, 160 minutes.
+That time might not be prohibitive, depending on the resources available; training is only needed when new student data is available.
+One major limitation in the results is that we cannot report AUROC results for yes/no criteria labels in the properties domain in Fig.
+3.
+In our dataset, class imbalance is extremely large for this domain, and no positive examples were included in the test set, making AUROC undefined.
+More data, especially more positive examples, are necessary for improvement.
+| AUROC (95% CI) | Train time | |
+| BoW | 0.839 (0.803, 0.873) | 1 min |
+| BERTfrozen | 0.887 (0.860, 0.913) | 7 min |
+| BERTfinetuned | 0.916 (0.893, 0.938) | 160 min |
+| Avg. (95% CI) | |
+| 0.029 (0.009, 0.049) | |
+| 0.077 (0.052, 0.104) | |
+| 0.048 (0.021, 0.075) |
+V Benefits and Limitations
+V.1 Utility.
+As we described in the introduction, we are working to automate the measurement of mechanistic sensemaking evidence for research purposes across a set of student responses to conceptual problems in introductory Newtonian mechanics.
+This will allow larger-scale research studies than is possible with human annotating alone.
+Currently, annotating problems by hand is burdensome.
+Our annotator required about 5 hours to apply our measurement scheme to student responses for one problem.
+In contrast, once trained the processing time for our models to annotate the same amount of data is less than 1 minute.
+Furthermore, training a model to perform this task is doable in less than 3 hours.
+We hypothesize that with much more labeled training data across many additional problems, the model proposed here could perform well on unseen problem instances in introductory mechanics.
 However, the amount of data needed to cross such a threshold is unclear.
-Regardless, initial efforts to address this limitation should be directed towards collaborative annotation.
-Reproducibility.
-The automation scheme ensures that the same student explanation receives the same annotation.
-The process is fully deterministic given fixed seeds to control for randomness.
-Randomness is only present in the initialization of the classifier weights and biases.
-Sustainability.
+Regardless, initial efforts to improve generalization across problems while reducing required annotation time should be directed towards collaborative annotation as well as semi-supervised learning methods [53].
+Training models to measure mechanistic sensemaking on more data from new problems opens the door for interesting studies that have been difficult for PER.
+Consider this example hypothesis: “Evidence of mechanistic sensemaking early in a course is better than early canonical correctness at predicting both sensemaking and correct understanding at the end of the course.” Testing that hypothesis would involve many hours of human analysis; the ML we have begun to develop would be able to carry out this investigation quickly and efficiently.
+ML-enabled measurement could be used to conduct a randomized controlled experiment in PER, such as the one in Kuo et al.
+[54], but specifically for understanding the variables that lead to more mechanistic sensemaking in students.
+It could also be used to evaluate various introductory physics problems as being of low or high quality depending on whether a large enough body of students displayed (or did not display) evidence of mechanistic sensemaking in their responses (e.g.
+problem 1 in our dataset).
+V.2 Generalizability
+Russ et al.
+[22] developed a framework for assessing mechanistic reasoning that has been in use in PER for almost 20 years.
+To develop ML, we adapt the categories based on a particular dataset to produce a scheme with 7 domains.
+Researchers with similar datasets within the same conceptual domain of classical mechanics may find the scheme and the ML it supports useful as is.
+New contexts, however, may require new projections of Russ’s et al’s framework.
+For example, evidence of mechanistic reasoning in student explanations of electricity and magnetism or quantum mechanics will look different.
+It is likely that analogies would be an essential domain for some subject matter.
+V.3 Reliability
+By designing and defining the 7 domains for our dataset, and providing annotated examples, we hoped to reduce ambiguity for human annotators - those who might look to reproduce results on our data or to apply the scheme to their own.
+To measure this, we conducted an inter-rater reliability (IRR) test.
+Two annotators who took smaller roles in development of the domains, one professor of computer science and one senior undergraduate student in computer science, used the scheme to measure the last 15 student responses across the four physics problems (60 student responses total).
+The first author provided them with a codebook containing: the domain definitions, the four problems, the criteria vectors for each problem, and 5 annotated student examples per problem with some explicit reasons as to why certain coding decisions were made.
+Annotators provided initial ratings, then engaged in a collaborative discussion to resolve ambiguity and clarify disagreement.
+After this discussion, new ratings were assessed.
+Among the three annotators (2 new raters alongside the first author who conducted all manual annotation for this work), the average accuracy and Cohen’s Kappa values [55] achieved pre-discussion are displayed in Table 3.
+The values that emerged post a discussion are presented in Table 3.
+| Ave. Accuracy | Ave. Cohen’s Kappa | |
+| Problem 1 | 1.0 | 1.0 |
+| Problem 2 | 0.86 | 0.66 |
+| Problem 3 | 0.82 | 0.63 |
+| Problem 4 | 0.83 | 0.48 |
+| Ave. Accuracy | Ave. Cohen’s Kappa | |
+| Problem 1 | 1.0 | 1.0 |
+| Problem 2 | 0.90 | 0.79 |
+| Problem 3 | 0.94 | 0.88 |
+| Problem 4 | 0.88 | 0.69 |
+Problem 1 showed perfect agreement, but very few of the student examples show evidence of mechanistic sensemaking.
+(This suggests a form of utility for instruction: perhaps this is not a useful problem to assign.) For the remaining three problems, IRR scores were reasonably high ( accuracy).
+Cohen’s Kappa values incorporate chance due to data imbalance and are lower across problems, showing moderate to substantial agreement according to the categories in [55].
+Discussion improved scores to accuracy, and all Cohen’s Kappa values increased to indicate substantial and almost perfect agreement [55].
+After discussion, our codebook language was updated to improve its clarity.
+For all results in Sec.
+IV, we trained ML to reproduce the labels provided by our single primary annotator.
+The 60 annotations from 2 other raters studied here were not included in our ML model training, as the IRR procedure was conducted in response to helpful reviewer feedback on the draft post model training based on the first author’s coding alone.
+As such, we highlight that this is a limitation of the current work - we do not know how our ML model will perform on measurements from multiple annotators nor do we have a model that corrects for mis-labeling.
+Investigating these directions is important for future larger-scale studies.
+Once an ML model is trained, the same student response will always receive the same annotation.
+Prediction is fully deterministic.
+V.4 Sustainability Privacy.
 As ML research advances, new encoders may be able to improve the overall classification performance.
@@
 As language modeling in ML improves, it is worth benchmarking new methods while also taking into account the amount of compute resources required to obtain only modestly better performance.
-III.3 ML for Education Related Works
-Machine learning tools are slowly integrating into STEM Ed.
-research methods, mostly with the introduction of text classification models [52, 43, 44, 45, 53].
-One focus is on in-context learning, where student text responses are directly fed into a chatbot window of a large language model to be labeled [42].
-This procedure can be conducted with a few examples (or zero) of student text, opening up an opportunity for non-ML practitioners to seamlessly use computation for data annotation.
-However, this procedure relies on the assumption that pre-trained language models are useful for specific tasks in education (e.g.
-classifying correctness, reasoning, sensemaking, etc.).
-Other works focus on training supervised ML models from scratch for their specific task [44, 45, 41].
-For example, Watts et al.
-[41] uses the mechanistic reasoning approach from Russ et al.
-to analyze student explanations in organic chemistry with the aid of ML binary classifiers.
-This work determines convolutional neural networks to outperform other model architectures for this task, but does not provide comparison results to other models.
-In our work, we provide a model comparison with three different language encoders (BoW, BERTfrozen, and BERTfinetuned), where we also fine-tune a pre-trained large language model on data for our specific task.
-Furthermore, we use a multi-label scheme that allows us to take advantage of shared structure across problems by providing each classifier with additional context for the problem-specific criteria we care about.
-As such, our classifiers are able to train on and classify new data across different physics problems without introducing new classification heads.
-In response to the increasing amount of supervised ML models in STEM Ed.
-research, Fussell et al.
-[54] effectively summarizes and provides a demonstration of the best practices widely used by ML practitioners for validating these ML models.
-We note that our work utilizes methods advised by this work, including both cross-validation and bootstrapping.
-Lastly, we note that while our work does not fall under the category of unsupervised learning, these models are a promising avenue of research in STEM Ed.
-and can be used to cluster student data, detect static latent variables in student text, and analyze the latent dynamics within student conversations [55].
-IV Practical Implementation Results
-IV.1 ML Model Deployability Analysis
-We seek to determine the deployability of the proposed classification models outlined in Sec.
-III.1.
-More specifically, our intention is to estimate how these models will perform given explanations from new students that it has never seen before, but on the same set of problems.
-Below, we outline our implementation and evaluation procedures prior to showcasing final results.
-Dataset of student explanations.
-Our dataset consists of 385 total de-identified student explanations across four problems distributed to students in the first few lectures of the 2023 Tufts University Introduction to Physics course.
-The retrospective analysis of this deidentified human subjects data was approved by our institutional review board.
-All problems have the same structure as Problem 4 II.1 shown above and are presented to the students as pre-lecture digital assignments [38].
-The same students are asked to solve each problem and explain how they obtained their solution.
-A few students chose not to respond to a particular problem, which brings our total number of explanations to .
-The problems capture different general core material for students to sensemake: (1) the descriptors of an object’s motion and dynamics; (2) the model of an object’s motion and dynamics under Earth’s gravitational influence; (3) the model of the relative motion of objects; (4) the model of an object’s motion dynamics under a net external influence.
-These problems vary in criteria vector size: (see Appendix A for exact problems and their details from [38]).
-The annotation procedure was carried out by a domain expert: the first author of this work who holds a doctorate in physics.
-Model and training specifics.
-In our analysis, we use the following models BoW, BERTfrozen, and BERTfinetuned.
-To evaluate performance, we randomly select 15 students who responded to all four questions to use as a held-out test set and use the rest of the data for training and validation (325 samples).
-To prevent data leakage, we stratify the data by students to ensure that responses from the same student do not end up in both training and test/validation sets.
-All models are trained with the Adam optimizer [56] for epochs.
-We perform a grid search over key model hyperparameters: learning rate and regularization strength ( for classifier weights and for BERT model weights), the full-details of which are available in Appendix D.
-For BoW and BERTfrozen, we use a 5-fold cross-validation approach for hyperparameter selection, where we select the hyperparameter combination with the best final average AUROC score across folds.
-AUROC is the computed area under a Receiver Operating Characteristic (ROC) curve defined by the true positive rate vs.
-the false positive rate of the predicted labels at different classification thresholds.
-There are two main reasons we believe AUROC to be the most useful metric for our context: (1) It does not require us to choose an arbitrary probability threshold to distinguish our positive predictions (label 1) from our negative predictions (label 0) for each classifier.
-Instead AUROC sums over all possible thresholds, providing a good summary that is still bounded, AUROC , with increasing values indicating better performance.
-(2) Due to imbalanced annotations (i.e.
-some classifiers have more of a positive or negative class rather than a 50-50 split) in our training dataset, AUROC is a safer metric to report than precision or accuracy, as the chance baseline is independent of imbalance.
-We include an annotation summary of the data in the Appendix B.
-The best model for BoW achieved an average AUROC score of across validation folds with the hyper-parameters: .
-Similarly, for BERTfrozen, the best model achieved an AUROC score of with the hyper-parameters: .
-For BERTfinetuned, we require a larger amount of computational resources per experiment, so we only used a single training and validation set to run the hyperparameter grid search.
-Note that we use an L2SP loss [57, 58] when training BERTfinetuned, consisting of two separate weight decay terms such that we can regularize the weights from the BERT encoder separately from the weights of our LR classifiers.
-This is done to best preserve the prior information learned in in BERT’s original training.
-We set the batch-size to , which has been shown empirically to work well with deep neural networks [59].
-Overall, we run 324 fine-tuning experiments for BERTfinetuned on Tufts’ HPC services.
-We picked the model with the optimal final AUROC of on the validation set, which had the following hyperparameters: , , and .
-Model evaluations.
-We retrain our selected models on all training and validation data, and evaluate on the held-out test set of randomly selected students (60 student explanations).
-To approximate uncertainty on unseen student explanations, we bootstrap our test set with replacement to generate individual test sets for evaluating performance [60, 61].
-In Figure 3, we show the average AUROC across independent test predictions for each of the 8 domain classifiers and for each model type.
-We are not able to report the AUROC for the properties domain as the class imbalance is extremely large and no positive examples were included in the test set.
-As such, we also provide the Area Under the Precision-Recall curve (AUPRC) results in Appendix E for all domain classifiers.
-This metric emphasizes the precision of the classifier, which informs us as to how well the model predicts the true positive labels.
-We note that even with AUPRC, there is too much uncertainty to make concrete claims about the deployability of the properties domain classifier.
-More positive data examples are necessary for improvement.
-We show the average AUROC across independent test set predictions for each model in Table 1.
-These values represent the annotation quality of our entire model.
-In Table 1, we also report the time taken to train the final selected model.
-To show the significance in results between models, we provide the confidence interval in Table 1 that results from our bootstrapping procedure and taking the AUROC differences between models.
-These values can be interpreted as the uncertainty intervals in AUROC difference across models, such that as long as zero is outside of the interval, we can be confident that there exists some non-zero difference between our models indicating that one is indeed better than the other.
-Additionally, we report the confusion matrices in Appendix F.
-BoW performs surprisingly well with an AUROC score of and trains efficiently in 1 minute.
-It performs nearly as well for objects as the other two models.
-This behavior is unsurprising as BoW reduces text to word frequency counts.
-BERTfrozen performs slightly better with an AUROC score of and takes 7 minutes to train.
-It performs the best for two of the domains: objects and mechanistic relationships.
-It is able to reliably outperform BoW on average, as 0 in not inside of the confidence interval between the AUROC classifier difference reported in Table 1.
-BERTfinetuned achieves an AUROC score of .
-It takes 160 minutes to train, but we emphasize that this is not a prohibitive cost for a model which only needs to be trained when new student data is available.
-It performs the best for influences, positioning, movements, interactions, and descriptive relationships.
-It is able to perform the best overall despite only leading in 5 domains due to the imbalance in the occurrence of each domain.
-We recommend the BERTfinetuned model as we believe the average increase in performance across domains is worth the nearly 3 hour training time, and the confidence intervals in 1 indicate it is a significant improvement over BERTfrozen.
-However, if the annotator wants to optimize for a particular domain, they should refer to Figure 3 to decide the best model for their purpose.
-| AUROC (95% CI) | Runtime | |
-| BoW | 0.839 (0.803, 0.873) | 1 min |
-| BERTfrozen | 0.890 (0.864, 0.915) | 7 min |
-| BERTfinetuned | 0.916 (0.893, 0.939) | 160 min |
-| Avg. (95% CI) | |
-| 0.026 (0.005, 0.045) | |
-| 0.077 (0.051, 0.104) | |
-| 0.051 (0.024, 0.079) |
-IV.2 Student Performance Analysis
-We provide a quantitative snapshot of student performance in the first few weeks of the 2023 Tufts Intro.
-to Physics course.
-More specifically, we compute the sensemaking scores of students from their explanations to four problems (used in Sec.
-IV.1 and displayed in Appendix Sec.
-A), and compare these scores alongside problem-solving accuracies.
-Students reported their multiple-choice answer, such that accuracy here is a binary variable .
-Our results in Fig 4 show that assessing sensemaking scores provides a more nuanced view of student performance than correctness alone.
-In problem 1, we observe a high proportion of students solving the problem correctly () with very low sensemaking scores among correct answers ().
-In problem 2, we also observe high correctness (), but with higher sensemaking scores (an average of ) among correct answers.
-In these problems we observe students who obtain a correct answer with both high and low evidence of sensemaking.
-In fact, a closer look at the data reveals that many of these students who arrived at the correct answer without showing evidence of physical sensemaking deployed a learned mathematical rule (i.e.
-the derivative of velocity is acceleration).
-In problems 3 and 4, we observe less correct answers across students ( and respectively); however sensemaking scores remain high among both correct (average sensemaking scores of and respectively) and incorrect ( and respectively) responses.
-Specifically in problem 4, the mean sensemaking scores for the incorrect group are higher than for the correct group.
-Recall that our measurement scheme determines evidence of the sensemaking process, not whether students have accurate information in the sensemaking process.
-Also, sensemaking has been characterized as a non-linear process to organize information that includes integrating ones phenomenological experiences, which often leads to constructive confusion [13, 14, 15, 16, 17].
-Investigating the student responses reveals that students with higher sensemaking scores who arrived at the incorrect answer either used incorrect information or were unable to successfully make connections to their personal experiences with everyday physical objects.
-| Problem | CI | |
-| 1 | -0.166 | (-0.461, 0.145) |
-| 2 | 0.053 | (-0.055, 0.176) |
-| 3 | 0.404 | (0.227, 0.546) |
-| 4 | -0.198 | (-0.391, 0.009) |
-We show the Point-Biserial correlation coefficient , known as the binary-continuous form of the Pearson correlation coefficient ), for the student data in Table 2.
-The coefficient , where values farther from zero signals a stronger linear relationship and the sign indicates the direction.
-Only results in problem 3 suggest that there exists a positive linear correlation between problem-solving accuracy and sensemaking score; however, the coefficient is closer to zero than one indicating that the relationship strength is weak ().
-As we cannot assume normality in the underlying distributions over sensemaking scores for each binary group, we circumvent a hypothesis test with a bootstrapping procedure.
-We sample our data with replacement times and generate an empirical distribution over scores for each sample.
-We report the confidence interval in Table 2.
-All values within each interval remain closer to zero than , except for problem 3 for the tail end of the confidence interval.
-The weak correlation present makes a clear point: problem-solving accuracy is not a useful proxy for measuring evidence of physical sensemaking, nor vice versa.
-If one were to only look at the accuracy scores, one might believe that students are learning the core material.
-Sensemaking scores add more nuance to this belief, suggesting that students are engaging in alternative methods to obtaining a problem solution.
-There are multiple actions one could take to further analyze these measurements and use them in a practical context, especially at the individual student level.
-For example, measurements could be used to observe whether evidence of sensemaking for a particular student is declining, stable, or growing across problems in time.
-To show an example of this, we show the trend for student in Fig.
-4.
-This student shows evidence of stable sensemaking growth, which does not always coincide with correctness.
-In fact, if one only observed the accuracy score, the student’s performance would appear to be declining.
-When both of these measures are combined, one observes a more nuanced picture of the student’s progress, which could lead to more informed interventions.
-V Conclusion & Future Work
-In this work, we contribute a post-hoc measurement scheme to quantify evidence of physical sensemaking via human annotation of student explanations to physics problems.
-To streamline the lengthy process of annotating each student explanation, we automate the measurement scheme with a trained ML model using language encoders and shared probabilistic classifiers.
-We implement three model variants with different language encoders and conduct a deployability analysis on real student explanations from Tufts’ 2023 Intro.
-to Physics course.
-We show that the ML models have good average performance on data from new students (0.916 AUROC for the best model), but performance is dramatically different across domains.
-Moreover, we show that across problems, we can now systematically measure evidence of student learning separately from problem-solving accuracy.
-In fact, we observe no linear correlation between these two variables.
-Such a result has important implications for both practical instruction and STEM Ed.
-research.
-First and foremost it suggests that we need to be more cautious with how we measure learning.
-It’s not always useful to make assumptions about the process from the outcome.
-Students obtaining correct answers does not necessarily indicate a student has engaged in physical sensemaking, as there are alternative strategies, oftentimes requiring less time in the short-term, to achieve the same goal.
-As previous works support, sensemaking can more directly be measured by paying attention to how students solve problems - whether that’s through a post-hoc written explanation or an in-moment oral discussion.
-The former is where we’ve focused our efforts, but we believe the latter can also be more efficiently executed with ML tools.
-Beyond this, we perceive two next-stage research questions that lie at the intersection of ML and STEM Ed.
-fields:
-(1) How can this post-hoc measurement scheme be improved?
-On the ML side, a natural first reaction is more data.
-We hypothesize that with more labeled training data across new problems, the models proposed here will be able to achieve better performance on the current task and generalize to unseen problems.
+We only consider models which are open-access, which allows us to train classifiers for our specific task and does not require us to share student data with a 3rd party.
+VI Conclusion & Future Work
+In this work, we contribute a PER-ML measurement scheme to quantify evidence of mechanistic sensemaking in student responses to introductory mechanics problems.
+By introducing an ML implementation, our scheme can be more efficiently applied to many student responses on many problems than manual coding.
+This can enable large-scale quantitative research studies in PER, though this technology has both promises and limitations (see Sec.
+V) and there is no “magic bullet.”
+In our work to date, we studied three model variants with different language encoders that require different levels of computational resources.
+We conducted a comparative analysis on real student responses.
+Results show that the ML models have good average performance on data from new students (0.916 AUROC for the best model).
+Performance does differ across domains of the scheme, in part reflecting the variance in the amount of available data per domain.
+Throughout the research process, we identified new open questions and challenges. We thus conclude with reflections that we hope will motivate further research.
+Recall we describe two steps of refinement from the prior framework of mechanistic reasoning [22].
+The first was to design a scheme for measuring evidence of mechanistic reasoning in the context of student explanations in introductory mechanics.
+The second step was to train ML based on human annotated measurements for 100 students across four mechanics problems.
+Reflections below consider these steps in reverse order.
+(1) How can the ML implementation of this measurement scheme be improved?
+A natural first reaction to this question is more data.
+We hypothesize that with more labeled training data across new problems, the supervised learning models proposed here will be able to achieve better performance on the current task and generalize to unseen problems.
 Given collaborative annotation efforts, such a model could be realized more efficiently and would save all users time in the long run.
-We invite interested collaborators to get in touch and encourage open sharing of sensemaking data when possible and appropriate.
-One could also experiment with new state-of-the-art encoders to enhance overall performance [62].
-For example, using the Mistral model (open access, open source transformer model in Ref.
-[63]) as the LLM text encoder, pre-trained with 7 billion parameters and 2 million language tokens.
-As off-the-shelf language models improve, one might not need as much training data to obtain quality performance.
-Lastly, our ML model assumes domain independence in the classification task.
-Perhaps there is a shared structure across domains that could be exploited.
+This also means that our model performance will depend on strong inter-rater agreement and accounting for human error in labeling.
+Considering future methods to account for different types of label noise is an interesting direction for future work.
+There are a few paths that we could take to improve performance.
+The first is to experiment with newer encoders from the rapidly-advancing state-of-the-art frontier [57].
+We could use the LLaMa encoder [2] previously used by Fussell et al.
+[8].
+For an even newer encoder, we could explore the open-access, open-source 7-billion parameter Mistral model [35].
+As off-the-shelf language models improve, one might not need as much task-specific training data to obtain quality performance, though it is hard to imagine doing well with less than our 4-problem, 100 student dataset.
+Task-specific training or in-context learning [58] will always be required.
+Out-of-the-box language encoders or encoder-decoder models (e.g.
+ChatGPT) are trained on general corpora of language, whereas the mechanistic sense-making of interest in PER reflects more particular disciplinary practices.
+Another approach to improve the ML would be to leverage the dependency structure across our domains.
 As a brief intuitive example, consider the perspective of an annotator.
-Given that the annotator identifies criteria in the first two domains (objects and influences), it is more probable that the annotator will identify criteria in the other domains.
-On the STEM Ed.
-side, we need to better understand how we can design problems that are aligned with the measurement scheme without creating a new performance metric for students.
-We do not want the annotation criterion to be perceived by students as tasks to complete.
-This runs the risk of students inventing alternative strategies to effectively mimic sensemaking.
-An example of this would be pattern recognizing the types of explanations that lead to high sensemaking scores.
-As shown by our ML model, this isn’t a very difficult task.
-Ideally, we prompt students to sensemake and our measurement scheme does not disturb their epistemic framing.
-There exists an open research pursuit in STEM Ed.
-to understand how problems can be effectively designed and communicated to students such that genuine sensemaking is encouraged [40].
-(2) How can we model the dynamics that lead to low vs.
-high evidence of physical sensemaking in the classroom?
-Our proposed measurement scheme only provides evidence of a student’s sensemaking at a single point in time.
-We lack a model as to how students evolve such that specific sensemaking scores are realized.
-It is of high interest to construct both deterministic and probabilistic ML mechanistic models that capture short and long-term developmental dynamics of student sensemaking in physics classrooms[64, 65].
+If the annotator identifies that the student response did not discuss even the central object in the objects domain, it is essentially almost certain that properties, spatial organizations, and movement activities are also not provided.
+This correlation between domains could be exploited for better modeling, and may provide valuable inductive bias for achieving quality performance with less labeled data.
+Lastly, we can pursue semi-supervised learning methods [53], which optimize model outputs to match labels that we do have, while also learning from patterns in a large, easier-to-collect un-labeled dataset (in our case, student response text without criteria vector annotations).
+We believe each of these paths forward are potentially valuable directions for improvement.
+(2) How can new useful measurement schemes be achieved?
+We set out, and we remain interested, to assess student reasoning for sensemaking, in particular sensemaking relevant to the disciplinary practices of physics.
+For the present effort, we limited our attention to mechanistic sensemaking, drawing on prior work in PER and working to connect it to progress in ML.
+We limited our attention further to introductory mechanics.
+Other areas such as electricity and magnetism or quantum mechanics, would need different measurement schemes for evidence of mechanistic sensemaking.
+We are not prepared to hypothesize that ML could create such schemes, but it would be worth exploration.
+There are also many other approaches to explore for incorporating insight from PER and ML research.
+There are other bases in the literature regarding evidence of sensemaking, broader than mechanistic.
+There are also other data modalities (e.g.
+hand-written solutions with diagrams, graphs, and sketches; audio and visual recordings).
+These choices will influence our ML decisions, and vice versa.
+Depending on the ML constraints (e.g.
+complexity and interpretability of the model, amount of data and compute resources needed), frameworks from the PER literature will need to be adapted.
+Thus, as shown in this work, we put forth that the research process that integrates PER and ML is interactive.
+Ideas from the PER and ML side will influence one another until convergence - which is why we want to encourage collaborations between longtime members of the PER community and ML researchers who are interested in bringing their domain expertise to advancing the mission of PER.
+It is really through these collaborations that we can develop meaningful schemes that take advantage of the efficiency promise of ML, and unlock new insights in understanding student thinking.
 Acknowledgements.
@@
 National Science Foundation via awards GCR and CAREER .
-Furthermore, we wish to acknowledge Prof.
-David Hammer for providing us with the dataset necessary to carry out this work, high-level feedback on the draft, and thoughtful discussions regarding STEM Ed.
-literature.
-Without him, this work would not have been possible.
-We wish to acknowledge other members of the Hughes lab research group, namely Ethan Harvey and Patrick Feeney, for thoughtful ML discussions.
+We wish to acknowledge other members of the Hughes lab research group, namely Ethan Harvey and Cynthia Feeney, for thoughtful ML discussions.
 Additionally, we’d like to thank the baristas at Jaho (Gabriel) and Starbucks (Ken) for supporting us through the final stages of this project.References
-- Adams and Wieman [2015] W. K. Adams and C. E. Wieman, Analyzing the many skills involved in solving complex physics problems, American Journal of Physics 83, 459 (2015).
-- Lestari et al.
-[2021] Lestari, S.
-Syafril, S.
-Latifah, E.
-Engkizar, D.
-Damri, Z.
-Asril, and N.
-E.
-Yaumas, Hybrid learning on problem-solving abiities in physics learning: A literature review, Journal of Physics: Conference Series 1796, 012021 (2021).
-- Docktor et al.
-[2016] J.
-L.
-Docktor, J.
-Dornfeld, E.
-Frodermann, K.
-Heller, L.
-Hsu, K.
-A.
-Jackson, A.
-Mason, Q.
-X.
-Ryan, and J.
-Yang, Assessing student written problem solutions: A problem-solving rubric with application to introductory physics, Phys.
+- Devlin et al. [2019] J. Devlin, M.-W. Chang, K. Lee, and K. Toutanova, Bert: Pre-training of deep bidirectional transformers for language understanding (2019).
+- Touvron et al.
+[2023] H.
+Touvron, T.
+Lavril, G.
+Izacard, X.
+Martinet, M.-A.
+Lachaux, T.
+Lacroix, B.
+Rozière, N.
+Goyal, E.
+Hambro, F.
+Azhar, A.
+Rodriguez, A.
+Joulin, E.
+Grave, and G.
+Lample, Llama: Open and efficient foundation language models, arXiv preprint arXiv:2302.13971 (2023).
+- BehnamGhader et al. [2024] P. BehnamGhader, V. Adlakha, M. Mosbach, D. Bahdanau, N. Chapados, and S. Reddy, Llm2vec: Large language models are secretly powerful text encoders (2024).
+- Haab et al. [2022] J. Haab, N. Deutschmann, and M. R. Martínez, Is attention interpretation? a quantitative assessment on sets (2022), arXiv:2207.13018 [cs.LG] .
+- Mylonas et al.
+[2022] N.
+Mylonas, I.
+Mollas, and G.
+Tsoumakas, An attention matrix for every decision: Faithfulness-based arbitration among multiple attention-based interpretations of transformers in text classification (2022), arXiv:2209.10876 [cs.CL] .
+- Wulff et al.
+[2025] P.
+Wulff, M.
+Kubsch, and C.
+Krist, eds., Applying Machine Learning in Science Education Research: When, How, and Why?, 1st ed., Springer Texts in Education, Vol.
+1 (Springer Cham, Cham, 2025) pp.
+XIII + 369, 19 b/w illustrations, 38 colour illustrations.
+- Krist et al.
+[2025] C.
+Krist, M.
+Kubsch, and P.
+Wulff, Introduction, in Applying Machine Learning in Science Education Research: When, How, and Why?, edited by P.
+Wulff, M.
+Kubsch, and C.
+Krist (Springer Nature Switzerland, Cham, 2025) pp.
+1–12.
+- Fussell et al.
+[2025] R.
+K.
+Fussell, M.
+Flynn, A.
+Damle, M.
+F.
+J.
+Fox, and N.
+G.
+Holmes, Comparing large language models for supervised analysis of students’ lab notes, Phys.
 Rev.
@@
 Res.
-12, 010130 (2016).
-- Burkholder et al.
-[2020] E.
-W.
-Burkholder, J.
-K.
-Miles, T.
-J.
-Layden, K.
-D.
-Wang, A.
-V.
-Fritz, and C.
-E.
-Wieman, Template for teaching and assessment of problem solving in introductory physics, Phys.
-Rev.
-Phys.
-Educ.
-Res.
-16, 010123 (2020).
-- Morphew et al.
-[2018] J.
-W.
-Morphew, J.
-P.
-Mestre, H.-A.
-Kang, H.-H.
-Chang, and G.
-Fabry, Using computer adaptive testing to assess physics proficiency and improve exam performance in an introductory physics course, Phys.
-Rev.
-Phys.
-Educ.
-Res.
-14, 020110 (2018).
-- Lipnevich et al.
-[2020] A.
-A.
-Lipnevich, T.
-R.
-Guskey, D.
-M.
-Murano, and J.
-K.
-S.
-and, What do grades mean?
-variation in grading criteria in american college and university courses, Assessment in Education: Principles, Policy & Practice 27, 480 (2020).
-- Fakcharoenphol et al. [2011] W. Fakcharoenphol, E. Potter, and T. Stelzer, What students learn when studying physics practice exam problems, Phys. Rev. ST Phys. Educ. Res. 7, 010107 (2011).
-- Ben Ouahi et al.
-[2021] M.
-Ben Ouahi, M.
-Ait Hou, A.
-Bliya, T.
-Hassouni, and E.
-M.
-Al Ibrahmi, The effect of using computer simulation on students’ performance in teaching and learning physics: Are there any gender and area gaps?, Education Research International 2021, 6646017 (2021).
-- Didi ş Körhasan and Pospiech [2024] N.
-Didi ş Körhasan and G.
-Pospiech, Examination of students’ quantum physics knowledge construction with peers and metacognitive evaluations, Phys.
-Rev.
-Phys.
-Educ.
-Res.
-20, 020153 (2024).
-- Laukenmann et al.
-[2003] M.
-Laukenmann, M.
-Bleicher, S.
-Fuß, M.
-Gläser-Zikuda, P.
-Mayring, and C.
-von Rhöneck, An investigation of the influence of emotional factors on learning in physics instruction, International Journal of Science Education 25, 489 (2003).
-- Faridi et al.
-[2021] H.
-Faridi, N.
-Tuli, A.
-Mantri, G.
-Singh, and S.
-Gargrish, A framework utilizing augmented reality to improve critical thinking ability and learning gain of the students in physics, Computer Applications in Engineering Education 29, 258 (2021).
-- Deslauriers et al.
-[2019] L.
-Deslauriers, L.
-S.
-McCarty, K.
-Miller, K.
-Callaghan, and G.
-Kestin, Measuring actual learning versus feeling of learning in response to being actively engaged in the classroom, Proceedings of the National Academy of Sciences 116, 19251 (2019).
+21, 010128 (2025).
+- Fussell et al. [2024] R. K. Fussell, E. M. Stump, and N. G. Holmes, Method to assess the trustworthiness of machine coding at scale, Phys. Rev. Phys. Educ. Res. 20, 010113 (2024).
 - Odden and Russ [2019a] T. O. B. Odden and R. S. Russ, Defining sensemaking: Bringing clarity to a fragmented theoretical construct, Science Education 103, 187 (2019a).
@@
 - Odden [2021] T. O. B. Odden, How conceptual blends support sensemaking: A case study from introductory physics, Science Education 105, 989 (2021).
-- Kapon [2017] S. Kapon, Unpacking sensemaking, Science Education 101, 165 (2017).
 - Danielak et al. [2014] B. A. Danielak, A. Gupta, and A. Elby, Marginalized identities of sense-makers: Reframing engineering student retention, Journal of Engineering Education 103, 8 (2014).
@@
 - Ford [2012] M. J. Ford, A dialogic account of sense-making in scientific argumentation and reasoning, Cognition and Instruction 30, 207 (2012).
+- Kapon [2017] S. Kapon, Unpacking sensemaking, Science Education 101, 165 (2017).
 - Russ et al.
@@
 Mikeska, Recognizing mechanistic reasoning in student scientific inquiry: A framework for discourse analysis developed from philosophy of science, Science Education 92, 499 (2008).
-- Kaldaras and Wieman [2023a] L.
-Kaldaras and C.
-Wieman, Instructional model for teaching blended math-science sensemaking in undergraduate science, technology, engineering, and math courses using computer simulations, Phys.
-Rev.
-Phys.
-Educ.
-Res.
-19, 020136 (2023a).
-- Kaldaras and Wieman [2023b] L.
-Kaldaras and C.
-Wieman, Cognitive framework for blended mathematical sensemaking in science, Int.
-Journal of STEM Education https://doi.org/10.1186/s40594-023-00409-8 (2023b).
-- KUO et al. [2013] E. KUO, M. M. HULL, A. GUPTA, and A. ELBY, How students blend conceptual and formal mathematical reasoning in solving physics problems, Science Education 97, 32 (2013).
-- Bain et al.
-[2018] K.
-Bain, J.-M.
-G.
-Rodriguez, A.
-Moon, and M.
-H.
-Towns, The characterization of cognitive processes involved in chemical kinetics using a blended processing framework, Chem.
-Educ.
-Res.
-Pract.
-19, 617 (2018).
-- Bing and Redish [2007] T. J. Bing and E. F. Redish, The cognitive blending of mathematics and physics knowledge, AIP Conference Proceedings 883, 26 (2007).
-- Kuo et al. [2020] E. Kuo, M. M. Hull, A. Elby, and A. Gupta, Assessing mathematical sensemaking in physics through calculation-concept crossover, Phys. Rev. Phys. Educ. Res. 16, 020109 (2020).
-- Gifford and Finkelstein [2020] J. D. Gifford and N. D. Finkelstein, Categorical framework for mathematical sense making in physics, Phys. Rev. Phys. Educ. Res. 16, 020121 (2020).
-- Dreyfus et al. [2017] B. W. Dreyfus, A. Elby, A. Gupta, and E. R. Sohr, Mathematical sensemaking in quantum mechanics: An initial peek, Phys. Rev. Phys. Educ. Res. 13, 020141 (2017).
-- Sherin [2001] B. L. Sherin, How students understand physics equations, Cognition and Instruction 19, 479 (2001).
-- Devlin et al. [2019] J. Devlin, M.-W. Chang, K. Lee, and K. Toutanova, Bert: Pre-training of deep bidirectional transformers for language understanding (2019).
-- Note [1] Github repo: https://github.com/tufts-ml/AuSeM.git.
-- Thagard and Kroon [2008] P. Thagard and F. Kroon, Hot Thought: Mechanisms and Applications of Emotional Cognition, A Bradford book (MIT Press, 2008).
 - Selen and Stelzer [2016] G. G. M. Selen and T. Stelzer, Classical mechanics (FlipItPhysics for University Physics. Macmillan Learning., 2016).
-- Cronbach [1951] L. J. Cronbach, Coefficient alpha and the internal structure of tests., Psychometrik 16, 297–334 (1951).
-- Amanda Benedict-Chambers and Palincsar [2017] E.
-A.
-D.
-Amanda Benedict-Chambers, Sylvie M.
-Kademian and A.
-S.
-Palincsar, Guiding students towards sensemaking: teacher questions focused on integrating scientific practices with science content, International Journal of Science Education 39, 1977 (2017).
+- Ko and Luna [2024] M.-L.
+M.
+Ko and M.
+J.
+Luna, The glue that makes it “hang together”: A framework for identifying how metadiscourse facilitates uncertainty navigation during knowledge building discussions, Journal of Research in Science Teaching 61, 457 (2024), https://onlinelibrary.wiley.com/doi/pdf/10.1002/tea.21893 .
+- Phillips et al. [2017] A. M. Phillips, J. Watkins, and D. Hammer, Problematizing as a scientific endeavor, Phys. Rev. Phys. Educ. Res. 13, 020107 (2017).
+- Machamer et al. [2000] P. Machamer, L. Darden, and C. F. Craver, Thinking about mechanisms, Philosophy of Science 67, 1 (2000).
+- Watts et al.
+[2020] F.
+Watts, J.
+Schmidt-McCormack, C.
+Wilhelm, A.
+Karlin, A.
+Sattar, B.
+Thompson, A.
+Gere, and G.
+Shultz, What students write about when students write about mechanisms: analysis of features present in students’ written descriptions of an organic reaction mechanism, Chemistry Education Research and Practice 21 (2020).
 - Watts et al.
@@
 Shultz, Developing machine learning models for automated analysis of organic chemistry students’ written descriptions of organic reaction mechanisms, in Student Reasoning in Organic Chemistry (The Royal Society of Chemistry, 2022).
-- Long et al. [2024] Y. Long, H. Luo, and Y. Zhang, Evaluating large language models in analysing classroom dialogue., npj Sci. Learn. (2024).
+- Aiken et al.
+[2021] J.
+M.
+Aiken, R.
+De Bin, H.
+J.
+Lewandowski, and M.
+D.
+Caballero, Framework for evaluating statistical models in physics education research, Phys.
+Rev.
+Phys.
+Educ.
+Res.
+17, 020104 (2021).
 - Ullmann [2019] T. Ullmann, Automated analysis of reflection in writing: Validating machine learning approaches, Int J Artif Intell Educ 29, 217–257 (2019).
@@
 18, 010141 (2022).
-- Qader et al. [2019] W. Qader, M. M. Ameen, and B. Ahmed, An overview of bag of words;importance, implementation, applications, and challenges (2019) pp. 200–204.
-- Murphy [2022] K. P. Murphy, Probabilistic Machine Learning: An introduction (MIT Press, 2022).
-- Turner [2024] R. E. Turner, An introduction to transformers (2024).
-- Sun et al. [2020] C. Sun, X. Qiu, Y. Xu, and X. Huang, How to fine-tune bert for text classification? (2020).
-- Gibaja and Ventura [2014] E. Gibaja and S. Ventura, Multi-label learning: a review of the state of the art and ongoing research, WIREs Data Mining and Knowledge Discovery 4, 411 (2014).
-- Bender et al.
-[2021] E.
-M.
-Bender, T.
-Gebru, A.
-McMillan-Major, and S.
-Shmitchell, On the dangers of stochastic parrots: Can language models be too big?, in Proceedings of the 2021 ACM Conference on Fairness, Accountability, and Transparency, FAccT ’21 (Association for Computing Machinery, New York, NY, USA, 2021) p.
-610–623.
-- Aiken et al.
-[2021] J.
-M.
-Aiken, R.
-De Bin, H.
-J.
-Lewandowski, and M.
-D.
-Caballero, Framework for evaluating statistical models in physics education research, Phys.
-Rev.
-Phys.
-Educ.
-Res.
-17, 020104 (2021).
 - Tschisgale et al.
@@
 19, 020123 (2023).
-- Fussell et al. [2024] R. K. Fussell, E. M. Stump, and N. G. Holmes, Method to assess the trustworthiness of machine coding at scale, Phys. Rev. Phys. Educ. Res. 20, 010113 (2024).
-- Hall and Krist [2025] K.
-Hall and C.
-Krist, Unsupervised ml with text data, in Applying Machine Learning in Science Education Research: When, How, and Why?, edited by P.
-Wulff, M.
-Kubsch, and C.
-Krist (Springer Nature Switzerland, Cham, 2025) pp.
-281–312.
+- Odden et al.
+[2020] T.
+O.
+B.
+Odden, A.
+Marin, and M.
+D.
+Caballero, Thematic analysis of 18 years of physics education research conference proceedings using natural language processing, Phys.
+Rev.
+Phys.
+Educ.
+Res.
+16, 010142 (2020).
+- Jiang et al.
+[2023] A.
+Q.
+Jiang, A.
+Sablayrolles, A.
+Mensch, C.
+Bamford, D.
+S.
+Chaplot, D.
+de las Casas, F.
+Bressand, G.
+Lengyel, G.
+Lample, L.
+Saulnier, L.
+R.
+Lavaud, M.-A.
+Lachaux, P.
+Stock, T.
+L.
+Scao, T.
+Lavril, T.
+Wang, T.
+Lacroix, and W.
+E.
+Sayed, Mistral 7b (2023).
+- Long et al. [2024a] Y. Long, H. Luo, and Y. Zhang, Evaluating large language models in analysing classroom dialogue., npj Sci. Learn. (2024a).
+- Long et al.
+[2024b] Q.
+Long, Y.
+Wu, W.
+Wang, and S.
+J.
+Pan, Does In-Context Learning Really Learn?
+Rethinking How Large Language Models Respond and Solve Tasks via In-Context Learning, in First Conference on Language Modeling (2024).
+- Min et al.
+[2022] S.
+Min, X.
+Lyu, A.
+Holtzman, M.
+Artetxe, M.
+Lewis, H.
+Hajishirzi, and L.
+Zettlemoyer, Rethinking the Role of Demonstrations: What Makes In-Context Learning Work?, in Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing, edited by Y.
+Goldberg, Z.
+Kozareva, and Y.
+Zhang (Association for Computational Linguistics, Abu Dhabi, United Arab Emirates, 2022) pp.
+11048–11064.
+- Jiang et al. [2020] Z. Jiang, F. F. Xu, J. Araki, and G. Neubig, How Can We Know What Language Models Know?, Transactions of the Association for Computational Linguistics 8, 423 (2020).
+- Liu et al.
+[2022] J.
+Liu, D.
+Shen, Y.
+Zhang, B.
+Dolan, L.
+Carin, and W.
+Chen, What Makes Good In-Context Examples for GPT-3?, in Proceedings of Deep Learning Inside Out (DeeLIO 2022): The 3rd Workshop on Knowledge Extraction and Integration for Deep Learning Architectures, edited by E.
+Agirre, M.
+Apidianaki, and I.
+Vulić (Association for Computational Linguistics, Dublin, Ireland and Online, 2022) pp.
+100–114.
+- Liang et al.
+[2021] P.
+P.
+Liang, C.
+Wu, L.-P.
+Morency, and R.
+Salakhutdinov, Towards Understanding and Mitigating Social Biases in Language Models, in Proceedings of the 38th International Conference on Machine Learning (PMLR, 2021) pp.
+6565–6576.
+- Qader et al. [2019] W. Qader, M. M. Ameen, and B. Ahmed, An overview of bag of words;importance, implementation, applications, and challenges (2019) pp. 200–204.
+- Murphy [2022] K. P. Murphy, Probabilistic Machine Learning: An introduction (MIT Press, 2022).
+- Turner [2024] R. E. Turner, An introduction to transformers (2024).
+- Sun et al. [2020] C. Sun, X. Qiu, Y. Xu, and X. Huang, How to fine-tune bert for text classification? (2020).
+- Gibaja and Ventura [2014] E. Gibaja and S. Ventura, Multi-label learning: a review of the state of the art and ongoing research, WIREs Data Mining and Knowledge Discovery 4, 411 (2014).
 - Kingma and Ba [2015] D.
@@
 - Raschka [2020] S. Raschka, Model evaluation, model selection, and algorithm selection in machine learning (2020), arXiv:1811.12808 [cs.LG] .
+- van Engelen and Hoos [2020] J. E. van Engelen and H. H. Hoos, A survey on semi-supervised learning, Machine Learning 109, 373 (2020).
+- Kuo et al. [2020] E. Kuo, M. M. Hull, A. Elby, and A. Gupta, Assessing mathematical sensemaking in physics through calculation-concept crossover, Phys. Rev. Phys. Educ. Res. 16, 020109 (2020).
+- Cohen [1960] J. Cohen, A coefficient of agreement for nominal scales, Educational and psychological measurement 20, 37 (1960).
+- Bender et al.
+[2021] E.
+M.
+Bender, T.
+Gebru, A.
+McMillan-Major, and S.
+Shmitchell, On the dangers of stochastic parrots: Can language models be too big?, in Proceedings of the 2021 ACM Conference on Fairness, Accountability, and Transparency, FAccT ’21 (Association for Computing Machinery, New York, NY, USA, 2021) p.
+610–623.
 - Zhao et al.
@@
 Wen, A survey of large language models (2024).
-- Jiang et al.
-[2023] A.
-Q.
-Jiang, A.
-Sablayrolles, A.
-Mensch, C.
-Bamford, D.
-S.
-Chaplot, D.
-de las Casas, F.
-Bressand, G.
-Lengyel, G.
-Lample, L.
-Saulnier, L.
-R.
-Lavaud, M.-A.
-Lachaux, P.
-Stock, T.
-L.
-Scao, T.
-Lavril, T.
-Wang, T.
-Lacroix, and W.
-E.
-Sayed, Mistral 7b (2023).
-- Wojnowicz et al.
-[2024] M.
-T.
-Wojnowicz, K.
-Gili, P.
-Rath, E.
-Miller, J.
-Miller, C.
-Hancock, M.
-O’Donovan, S.
-Elkin-Frankston, T.
-T.
-Brunyé, and M.
-C.
-Hughes, Discovering group dynamics in coordinated time series via hierarchical recurrent switching-state models (2024).
-- Lerner et al.
-[2006] R.
+- Brown et al.
+[2020] T.
+B.
+Brown, B.
+Mann, N.
+Ryder, M.
+Subbiah, J.
+Kaplan, P.
+Dhariwal, A.
+Neelakantan, P.
+Shyam, G.
+Sastry, A.
+Askell, S.
+Agarwal, A.
+Herbert-Voss, G.
+Krueger, T.
+Henighan, R.
+Child, A.
+Ramesh, D.
 M.
-Lerner, W.
-Damon, E.
-Thelen, and L.
-B.
-Smith, Dynamic systems theories, in Handbook of Child Psychology.
-Volume 1, Theoretical Models of Human Development (John Wiley & Sons, 2006) 6th ed., pp.
-258–312.
-Appendix A Physics Problems & Annotations
+Ziegler, J.
+Wu, C.
+Winter, C.
+Hesse, M.
+Chen, E.
+Sigler, M.
+Litwin, S.
+Gray, B.
+Chess, J.
+Clark, C.
+Berner, S.
+McCandlish, A.
+Radford, I.
+Sutskever, and D.
+Amodei, Language models are few-shot learners (2020), arXiv:2005.14165 .
+Appendix A Physics problems with criteria questions
 We provide all four physics problems as they originally appeared to students from Ref.
@@
 Annotation Procedure 1
-Components
 (1) Does the student identify the object (object, car, etc.)?
-Characteristics
-(2) Does the student identify the orientation (e.g. left or right)?
-(3) Does the student identify movement 1 (e.g. slow down)?
-(4) Does the student identify movement 2 (e.g. zero)?
-(5) Does the student identify movement 3 (e.g. speed up)?
-Relationships
-(6) Does the student identify the descriptive relationship (e.g position changes is a pos/neg velocity)?
-(7) Does the student identify the descriptive relationship (e.g velocity changes is a pos/neg acceleration)?
+(2) Does the student identify the object orientation (e.g. left or right)?
+(3) Does the student identify movement 1 of the object(e.g. slow down)?
+(4) Does the student identify movement 2 of the object (e.g. zero)?
+(5) Does the student identify movement 3 of the object (e.g. speed up)?
 Problem 2
@@
 Annotation Procedure 2
-Components
 (1) Does the student identify object 1 (e.g. ball 1)?
@@
 (4) Does the student identify the possible environmental influences on object 2 (e.g. gravity)?
-Characteristics
-(5) Does the student identify object 1 positioning (e.g. down)?
-(6) Does the student identify object 2 positioning (e.g. horizontal up)?
-(7) Does the student identify object 1 motion (e.g. speeding up)?
-(8) Does the student identify object 2 motion (e.g. speeding up)?
-Relationships
+(5) Does the student identify object 1 spatial organization (e.g. down)?
+(6) Does the student identify object 2 spatial organization (e.g. horizontal up)?
+(7) Does the student identify object 1 motion post the drop (e.g. speeding up)?
+(8) Does the student identify object 2 motion post the drop (e.g. speeding up)?
 (9) Does the student identify the interaction between object 1 and gravity (e.g. gravitational force)?
@@
 Annotation Procedure 3
-Components
 (1) Does the student identify object 1 (e.g. girl)?
@@
 (3) Does the student identify object 3 (e.g. moving sidewalk, conveyor belt)?
-Characteristics
-(4) Does the student identify the positioning of the dog (e.g. left)?
-(5) Does the student identify the positioning of the belt (e.g. right)?
+(4) Does the student identify the spatial organization of the dog (e.g. left)?
+(5) Does the student identify the spatial organization of the belt (e.g. right)?
 (6) Does the student identify the movement of the girl (e.g. at rest)?
@@
 (8) Does the student identify the movement of the belt (e.g. 2 m/s constant speed)?
-Relationships
 (9) Does the student identify the interaction between the girl and the belt (e.g. moving her to the right)?
@@
 (e.g.
-the speed of the observer relative to the observed object in motion causes the observed motion)
+the observer moving relative to the object impacts the observed motion of the object)
 Problem 4
@@
 Annotation Procedure 4
-Components
 (1) Does the student identify the object (e.g. object, box, etc.)?
 (2) Does the student identify the possible environmental influence acting on the object (e.g. applied net force)?
-Characteristics
 (3) Does the student identify the object properties (e.g. mass)?
-(4) Does the student identify the object positioning before applied force (e.g. upwards)?
-(5) Does the student identify the object positioning after applied force (e.g. horizontal)?
+(4) Does the student identify the object spatial organization before applied force (e.g. upwards)?
+(5) Does the student identify the object spatial organization after applied force (e.g. horizontal)?
 (6) Does the student identify the object movement prior to the applied force (e.g. constant velocity)?
 (7) Does the student identify the object movement after the applied force (e.g. final velocity)?
-Relationships
 (8) Does the student identify the interaction between the object and the net applied force (e.g. change in direction)?
@@
 Appendix B Annotation tables
-We provide an annotation summary. We showcase the number of student explanations per problem, and the number of those explanations being labeled as positive for each specific criteria.
+We provide an annotation summary. We showcase the number of student responses per problem, and the number of those explanations being labeled as positive for each specific criteria.
 | Problem | Number of Responses |
@@
 | 5 | 4 | 40 | 9 | 46 |
-| 6 | 6 | 40 | 80 | 20 |
-| 7 | 63 | 46 | 92 | 44 |
+| 6 | 40 | 80 | 20 | |
+| 7 | 46 | 92 | 44 | |
 | 8 | 46 | 29 | 41 | |
@@
 Given the training dataset of assumed independent student embeddings with human annotated labels, we train all parameters to minimize the following loss function:
-Here, is the total number of problems, is the total number of student explanations per problem in the dataset , is the total number of domains, and is the set of all context criteria (indexed by ) in a particular domain for a specific problem .
-In the binary cross-entropy expression, is the ground truth prediction for the student explanation associated with problem with respect to domain and specific criteria .
+Here, is the total number of problems, is the total number of student responses per problem in the dataset , is the total number of domains, and is the set of all context criteria (indexed by ) in a particular domain for a specific problem .
+In the binary cross-entropy expression, is the ground truth prediction for the student response associated with problem with respect to domain and specific criteria .
 is the corresponding predictive probability.
+Note that by ground truth, we mean the human-annotated labels provided in this work.
 The loss representing fit to the training data is averaged over all examples.
@@
 Appendix E AUPRC Scores
-We show the AUPRC scores for each independent classifier in Fig. 5 and the average AUPRC scores across models in Table 5.
-| AUPRC (95% CI) | Runtime | |
-| BoW | 0.831 (0.788, 0.870) | 1 min |
-| BERTfrozen | 0.881 (0.846, 0.911) | 7 min |
-| BERTfinetuned | 0.896 (0.863, 0.927) | 160 min |
+We show the test set AUPRC scores across domains for each classifier in Fig. 4. The micro-averaged test set AUPRC score for each model is in Table 6.
+| AUPRC (95% CI) | Train time | |
+| BoW | 0.829 (0.786, 0.868) | 1 min |
+| BERTfrozen | 0.876 (0.842, 0.906) | 7 min |
+| BERTfinetuned | 0.895 (0.862, 0.926) | 160 min |
 Appendix F Confusion matrices
-We present confusion matrices for all three methods.
+Below, we present for each method a confusion matrix covering all yes/no criteria questions for the 15-student test set.
 To produce binary predictions, we threshold the predicted probability at a value of 0.5.
@@
 BERTfinetune has the fewest classification errors overall.
-However, we see that BERTfrozen makes slightly fewer false positive mistakes (45 compared to 47).
-For cases where one type of error has a larger cost than another, careful model and threshold selection are important.
+However, we see that BERTfrozen makes the same number of false positive mistakes (45).
+For deployment applications where one type of error has a larger cost than another, careful model and threshold selection are important.
+BoW on test set:
+BERTfrozen on test set:
+BERTfinetuned on test set: